### v1

# XXLTraffic: Expanding and Extremely Long Traffic Dataset for Ultra-Dynamic Forecasting Challenges

Abstract
Traffic forecasting is crucial for smart cities and intelligent transportation initiatives, where deep learning has made significant progress in modeling complex spatio-temporal patterns in recent years. However, current public datasets have limitations in reflecting the ultra-dynamic nature of real-world scenarios, characterized by continuously evolving infrastructures, varying temporal distributions, and temporal gaps due to sensor downtimes or changes in traffic patterns. These limitations inevitably restrict the practical applicability of existing traffic forecasting datasets. To bridge this gap, we present XXLTraffic, the largest available public traffic dataset with the longest timespan and increasing number of sensor nodes over the multiple years observed in the data, curated to support research in ultra-dynamic forecasting. Our benchmark includes both typical time-series forecasting settings with hourly and daily aggregated data and novel configurations that introduce gaps and down-sample the training size to better simulate practical constraints. We anticipate the new XXLTraffic will provide a fresh perspective for the time-series and traffic forecasting communities. It would also offer a robust platform for developing and evaluating models designed to tackle ultra-dynamic and extremely long forecasting problems. Our dataset supplements existing spatio-temporal data resources and leads to new research directions in this domain.
1 Introduction
Traffic prediction plays a vital role in intelligent transportation systems by offering accurate and dependable forecasts of future traffic conditions that serve as critical references for route design, urban construction, and road management. The long-term spatio-temporal relationship is essential and irreplaceable for traffic forecasting, i.e., the long-term relationship can not only provide context that helps understand anomalies or irregularities in short-term data. but also help identify patterns and trends over different timescales, such as daily, weekly, monthly, and yearly cycles. In recent years, significant work has focused on short-term and long-term traffic flow prediction. Various deep learning techniques, such as GNNs, emphasize extracting spatial relationships in traffic networks [13], while Transformer variants and upgraded architectures and methods focus on capturing short-term and long-term temporal relationships [29]. Existing work has achieved promising performance in both short-term and long-term predictions, but this is limited to relatively simple experimental setups.
The inevitable tendency in intelligent transportation systems is to design models that incorporate more realistic circumstances rather than just simple experimental conditions. This requires the use of data sets that accurately reflect real-world scenarios. This serves as a strong incentive for us to create and develop XXLTraffic, i.e., why does XXLTraffic focus on ultra-dynamic forecasting? Why is it necessary to include extremely long-term data? We will introduce the commonalities and developments of existing datasets, as well as the difficulties in establishing XXLTraffic more specifically.
Figure 1: Our dataset are evolving and longer than existing datasets. Existing datasets are either limited by short temporal spans or insufficient spatial nodes. In contrast, our dataset features an evolving growth of spatial nodes and spans over 20 years.
Recent Advances in Expanding Traffic Datasets. Real-world traffic scenarios necessitate more complex prediction settings, involving extended temporal horizons or broader spatial coverage in experiments. In the temporal domain, new settings are typically proposed based on previously published work rather than introducing new datasets:  [30] and [11] expanded input and output lengths up to four times on existing datasets.From a spatial perspective, [4] published a dataset with nodes growing annually and provided an evolving network to support new node predictions. [33] proposed a continual learning framework with pattern expansion mechanisms based on [4]. Additionally, [27] and [25] offered larger-scale spatial node datasets to support subsequent researchers. Recent work has explored longer temporal step experimental settings and released traffic datasets spanning up to five years and thousands of nodes. However, in specific scenarios, such as future traffic prediction for highway planning, these data and experimental settings fall short. As shown in Figure 1, most existing datasets have limitations in temporal span, which inspired us to develop a dataset for expanding and extremely long traffic forecasting. Highway construction typically spans several years, requiring predictions several years in advance. These extremely long prediction scenarios necessitate traffic datasets with even longer temporal spans for support. As the temporal span extends, urban infrastructure development and road construction can lead to shifts in traffic patterns, resulting in an evolving domain shift. This observation motivated us to provide an expanding and extremely long traffic dataset. Additionally, the combination of these factors enables the extraction of more temporal patterns from extremely long sequences, allowing for the possibility of longer input sequences.
Challenges. The ultra-dynamic challenge encompasses three key aspects: (1) Continuously evolving states of the underlying spatio-temporal infrastructures, characterized by an expanding number of nodes over the years. This continuous growth introduces complexity as the infrastructure adapts and expands. (2) Evolving temporal distributions over an extremely long observation period, which is crucial for extremely long forecasting. This requires models to adapt to changes in patterns and trends over extensive temporal spans. (3) Ultra spatio-temporal dynamicity with temporal gaps. These gaps can occur due to various reasons such as sensor data becoming unavailable during road construction, maintenance activities, or shifts in data distributions. This dynamic nature necessitates robust modeling techniques to handle inconsistencies and interruptions in data.
Contribution. We have constructed a traffic dataset with an exceptionally long temporal span and broader regional coverage, providing aggregated data and benchmarking, as well as a benchmarking setup considering extremely long prediction scenarios for future exploration:
  • We propose XXLTraffic, a dataset that spans up to 23 years and exhibits evolutionary growth. It includes data from 9 regions, with detailed data collection and processing procedures for expansion and transformation. This dataset supports both temporally scalable and spatially scalable challenges in traffic prediction.
  • We present an experimental setup with temporal gaps for extremely long prediction with gaps scenarios and provide a benchmark of aggregated versions of hourly and daily datasets.
  • We provide the exploration of input features through evolving temporal distributions over an extremely long observation period. Additionally, our datasets support zero-shot forecasting for new sensors.
2 Preliminaries
In this section, we will define traffic data and traffic prediction tasks.
Definition 1. Traffic datasets:
Traffic data primarily consists of vehicle flow detection data collected by sensors distributed across various locations in the traffic network. It is generally represented by $X_{i}$in $\mathbb{R}^{N\times T\times C}$, where $T$ denotes the time steps, $N$ denotes the number of sensors, and $C$ denotes the number of features.
Definition 2. Short-term traffic prediction:
Short-term traffic prediction primarily focuses on forecasting traffic speed or flow within the next hour. As shown in Equation 1, the input length $\alpha$ and output length $\beta$ are generally set to 12 steps.
$[X_{t-(\alpha-1)},...,X_{t-1},X_{t}]\rightarrow[X_{t+1},X_{t+2},...,X_{t+\beta}],$ (1)
Definition 3. Long-term multivariate prediction:
This task mainly focuses on long-sequence time series prediction, which includes the traffic dataset. As shown Table 1, the sequence length can reach up to 2880 steps.
Definition 4. Extremely Long Prediction with Gaps:
Based on Equation 1, the observation and prediction are not adjacent but are instead separated by a gap period $g$, as shown in the following formula.
$[X_{t-(\alpha-1)},...,X_{t-1},X_{t}]\rightarrow[X_{t+g+1},X_{t+g+2},...,X_{t+g+\beta}],$ (2)
3 Gaps and Comparison with Existing Traffic Datasets
Figure 2: Our expanding and extremely long prediction addresses the existing limitations in both short-term and long-term predictions.
As shown in Table 1, existing traffic prediction work can easily be divided into short-term and long-term settings. The short-term setting originated from the STGCN[41] work, while the long-term setting was first introduced by LSTNet [14] and subsequently established as a widely adopted experimental framework by Informer [45]. In recent years, short-term prediction typically has a maximum step length of 12 steps, while long-term prediction reaches up to 720 steps. However, works such as Witran [11] and DAN [19] recognized the need for even longer step predictions in practical applications, extending the length to a maximum of four times the typical length. Despite the differences in step lengths, their observed and predicted values are concatenated tightly together, as shown in Equation 1. To accommodate complex real-world scenarios, such as highway route planning predictions, it is necessary to introduce a gap of several years between observation and prediction. Typically, existing datasets lack the temporal coverage required to support gaps exceeding one year.At the same time, predicting several years in advance also implies the need to forecast traffic for sensors at new locations, taking into account the evolving nature of the road network. Even when such coverage is available, works like [33] and [4] utilize evolving datasets but do not provide sufficient data to train models for extended durations. To overcome these gaps, our Expanding and Extremely Long Traffic Dataset robustly supports these complex scenarios.
Table 1: Summary of recent short-term traffic forecasting and long-term multivariate forecasting
Datasets Model Series Length
Short-term STGCN [41] {3,6,9,12}
DCRNN [20], GWN [40], BTF [3], DMSTGCN [10], {3,6,12}
GTS [28], STGODE [7], PM-MemNet [16], STAEFormer [23]
AGCRN [1], STSGCN [32], ,DSTAGNN [15] D2STGNN [31], {12}
DyHSL [44],PDFormer [12], MultiSPANS [47], GMSDR [22]
Long-term MTGNN [39],LSTNet [14] {3,6,12,24}
ARU [6] {12,24,48,168,336}
LogSparse_Trans [18] {24,48,72,96,120,144,168,192}
AST [38] {8,24,168,336}
SSDNet [21] {20,24,30,138}
Informer [45], Autoformer [37], FEDformer [46], {24,48,96,192,336,720}
Linear [17], Triformer [5], Pyraformer [24]
DSformer [42],DeepTime [35],DLinear [43]
Witran [11] {168, 336, 720, 1440, 2880}
DAN [19] {288, 672, 1440}
4 The XXLTraffic Datasets
4.1 Data Collection
We obtained the expanding and extremely long traffic sensor data from the California Department of Transportation (CalTrans) Performance Measurement System [^1] https://pems.dot.ca.gov/ (PeMS) [2]. PeMS is an online platform that collects traffic data from 19,561 sensors distributed across California state highways. These sensor locations are divided into nine districts. We downloaded all the raw data for these nine districts from the initial data release up to March 20, 2024. The system automatically generates a daily data file for each district, containing data from all sensors within each district. We have stored the complete raw data files in an open-source repository for quick access, which will be released after the publication.
Table 2: Comparison between our XXLTraffic dataset and the existing traffic datasets.
Reference Dataset Samples Nodes Time Interval Time Span Time Period
DCRNN [20] METR-LA 34,272 207 5 mins 4 months 03/2012 - 06/2012
PEMS-BAY 52,116 325 5 mins 6 months 01/2017 - 05/2017
LSTNet [14] Traffic 17,544 862 1 hour 2 years 01/2015 - 12/2016
STGCN [41] PEMSD7(M) 12,672 228 5 mins 2 months 05/2012 - 06/2012
PEMSD7(L) 12,672 1026 5 mins 2 months 05/2012 - 06/2012
ASTGCN [9] PEMSD4-I 17,002 228 5 mins 2 months 01/2018 - 02/2018
PEMSD8-I 17,856 1,979 5 mins 2 months 07/2016 - 08/2016
STSGCN [32] PEMS03 26,208 358 5 mins 11 months 01/2018 - 11/2018
PEMS04 16,992 307 5 mins 2 months 01/2018 - 02/2018
PEMS07 28,224 883 5 mins 2 months 05/2017 - 06/2017
PEMS08 17,856 170 5 mins 2 months 07/2016 - 08/2016
Large-ST [25] CA 525,888 8,600 5 mins 5 years 01/2017 - 12/2021
GLA 525,888 3,834 5 mins 5 years 01/2017 - 12/2021
GBA 525,888 2,352 5 mins 5 years 01/2017 - 12/2021
SD 525,888 716 5 mins 5 years 01/2017 - 12/2021
Ours Full_PEMS03 2,419,488 1809 5 mins 23.00 years 03/2001 - 03/2024
Full_PEMS04 2,287,872 4,089 5 mins 21.75 years 06/2002 - 03/2024
Full_PEMS05 1,998,720 573 5 mins 19.00 years 03/2005 - 03/2024
Full_PEMS06 1,945,728 705 5 mins 18.50 years 09/2005 - 03/2024
Full_PEMS07 2,287,872 4,888 5 mins 21.75 years 06/2002 - 03/2024
Full_PEMS08 2,419,488 2,059 5 mins 23.00 years 03/2001 - 03/2024
Full_PEMS10 1,998,720 1,378 5 mins 19.00 years 03/2005 - 03/2024
Full_PEMS11 2,261,376 1,440 5 mins 21.50 years 09/2002 - 03/2024
Full_PEMS12 2,331,360 2,587 5 mins 22.16 years 01/2002 - 03/2024
As illustrated in Table 2, our collected dataset significantly exceeds existing datasets in terms of both temporal coverage and the number of spatial nodes. The dataset sample will be available on: https://github.com/cruiseresearchgroup/XXLTraffic, which includes the raw data, sensor metadata (containing sensor IDs, geographical coordinates, associated road information, etc.), the data processing pipeline code, and the processed datasets.
4.2 Data Preprocessing
Based on the 23 years of raw data we collected, we conducted rigorous data filtering and aggregation. The PeMS system has continuously evolved, expanding from a few sensors in 2001 to over 4,000 sensors in some districts today. To support our setting of extremely long forecasting with gaps, we selected a subset of sensors that were installed in the early stages and have consistently collected new data up to the present(named gap dataset), which is shown in the Appendix. This extensive gap dataset effectively underpins the extremely long forecasting with gaps demonstrated in Figure 2. Utilizing the gap dataset, we performed both hourly and daily aggregations, which will be employed for gap-free long-term forecasting benchmarking. We will provide standard long-term forecasting benchmarks for both the hourly and daily datasets.
4.3 Data Overview
The XXLTraffic dataset is distributed across highways in the state of California, as illustrated in the Figure 3(a). The nine colors represent nine districts. From Figures 3(b),  3(c), and 3(d), we can clearly observe the evolutionary growth of the sensors. The sensors are extensively distributed across both urban and suburban areas, offering diverse modalities. Additionally, the sensors are densely interconnected, enabling the formation of a high-quality traffic graph dataset.
(b) District 08 in 2005
(c) District 08 in 2015
(d) District 08 in 2024
Figure 3: XXLTraffic dataset overview and its evolving development. This figure provides a global overview and two local overviews, showcasing the diversity of sensor distribution. The lower parts highlights a selected region to illustrate the growth and changes in traffic sensors over time.
It is evident that sensors at the same location may collect completely different distributions over the course of urban evolution. As shown in Figure 4, some sensors have maintained the same distribution from 2005 to 2024, while others have experienced significant changes in distribution. The temporal changes causing domain shifts present a significant challenge for our extremely long forecasting.
Figure 4: Sensor traffic status distribution of District 8 from 2005 to 2024. While some sensors exhibit minimal changes, others show significant distribution differences, regardless of whether they are in low-traffic or high-traffic areas. This presents substantial challenges for extremely long forecasting with long gaps.
4.4 XXLTraffic Licence
The XXLTraffic dataset is licensed under CC BY-NC 4.0 International: https://creativecommons.org/licenses/by-nc/4.0. Our code is available under the MIT License: https://opensource.org/licenses/MIT. Please check the official repositories for the licenses of any specific baseline methods used in our codebase.
5 Experiments
We conducted experiments for both extremely long forecasting with gaps using gap dataset and conventional long-term forecasting using hourly dataset and daily dataset. Additionally, referring to the definition in Figure 2, we set the gap parameter $g$ as 1 year, 1.5 years, and 2 years for the gap dataset, as illustrated by Figure 5.
Figure 5: Problem definition. The yellow boxes represent typical predictions, the gray boxes denote gap periods between observation and prediction, and the blue boxes indicate extended predictions.
5.1 Datasets
We conducted experiments on all proposed sub-datasets. To maintain consistency with previous state-of-the-art benchmarks, we selected districts 03, 04, 07, and 08 (widely recognized as PEMS03/04/07/08) for our experiments, using the gap, hourly, and daily datasets from these districts. Results for other datasets are presented in Appendix. All sub-datasets were divided into training, validation, and test sets using a 6:2:2 ratio. For the gap dataset, due to the extensive span of up to 20 years resulting in a large sample size, we fixed a seed during data preprocessing to select 10% of the dataset for training and testing to quickly demonstrate our results. In our experiments, due to GPU memory constraints, we pruned PEMS07_ours and PEMS12_ours by removing nodes with a missing data ratio greater than 10%, reducing the node number to 1,613 and 867. The details of the datasets used in our benchmarking is in Appendix A.1.
Table 3: Comparison in gap dataset. The bold text indicates the best.
Methods Mamba iTransformer DLinear Autoformer
Metrics MSE MAE MSE MAE MSE MAE MSE MAE
PEMS03_gap 1-year gap 96 1.457 0.913 1.636 0.989 1.500 0.933 1.301 0.906
192 1.472 0.922 1.597 0.970 1.542 0.945 1.266 0.867
336 1.434 0.913 1.512 0.935 1.531 0.935 1.137 0.807
1.5-year gap 96 1.485 0.942 1.879 1.078 1.653 0.976 1.467 0.945
192 1.442 0.934 1.753 1.017 1.642 0.968 1.460 0.923
336 1.446 0.938 1.662 0.987 1.632 0.968 1.345 0.882
2-year gap 96 1.359 0.833 1.844 1.048 1.568 0.954 1.328 0.894
192 1.216 0.772 1.729 1.008 1.473 0.896 1.235 0.837
336 1.294 0.801 1.614 0.960 1.407 0.870 1.220 0.838
PEMS04_gap 1-year gap 96 1.325 0.914 1.644 1.037 1.396 0.955 0.819 0.717
192 1.438 0.642 1.587 1.018 1.440 0.965 0.941 0.767
336 1.424 0.935 1.447 0.960 1.421 0.954 0.853 0.730
1.5-year gap 96 1.204 0.890 2.014 1.193 1.488 0.995 0.981 0.779
192 0.982 0.787 1.649 1.038 1.301 0.913 0.867 0.735
336 0.961 0.792 1.352 0.929 1.298 0.908 0.762 0.679
2-year gap 96 1.220 0.893 1.652 1.074 1.446 0.977 0.909 0.755
192 1.004 0.807 1.189 0.870 1.268 0.893 0.909 0.747
336 1.198 0.864 1.584 1.016 1.269 0.898 0.898 0.736
PEMS07_gap 1-year gap 96 1.719 1.006 1.756 1.024 1.680 1.024 1.215 0.867
192 1.637 0.996 1.816 1.061 1.776 1.043 1.202 0.852
336 1.637 0.991 1.784 1.028 1.774 1.034 1.098 0.774
1.5-year gap 96 1.571 0.972 1.705 1.016 1.585 0.987 1.184 0.844
192 1.578 0.982 1.657 1.004 1.624 0.996 1.088 0.800
336 1.818 1.127 1.712 1.028 1.668 1.036 0.991 0.729
2-year gap 96 5.411 1.319 2.055 1.149 1.746 1.048 1.099 0.802
192 12.620 1.370 2.006 1.128 1.705 1.002 1.284 0.875
336 9.614 1.378 1.812 1.057 1.605 0.968 0.972 0.734
PEMS08_gap 1-year gap 96 5.411 1.319 1.514 0.950 1.771 1.058 1.153 0.843
192 12.620 1.370 1.499 0.900 1.762 1.059 1.202 0.845
336 9.614 1.378 1.654 0.984 1.961 1.131 1.184 0.849
1.5-year gap 96 4.453 1.061 2.286 1.196 1.978 1.072 1.428 0.901
192 9.413 1.046 1.713 0.971 1.606 0.928 1.314 0.833
336 10.457 1.063 1.890 1.488 1.736 0.985 1.320 0.836
2-year gap 96 5.117 0.953 2.400 1.232 1.969 1.069 1.474 0.895
192 12.769 0.933 1.974 1.070 1.670 0.942 1.505 0.883
336 13.382 0.946 1.936 1.056 1.628 0.926 1.464 0.874
5.2 Baselines
In our comparison experiments, we adopted four popular baselines, including MLP, Transformer, and Mamba architectures. Autoformer [37], an earlier state-of-the-art model, leverages a decomposition architecture and auto-correlation mechanism to enhance efficiency and accuracy in long-term time series forecasting, outperforming traditional Transformer models. iTransformer [26] is the latest and most effective Transformer-based model, utilizing attention and feed-forward networks on inverted dimensions, embedding time points into variate tokens. DLinear [43] challenges the effectiveness of Transformer models by proposing a simple one-layer linear model that captures temporal relations in an ordered set of continuous points. It employs positional encoding and uses tokens to embed sub-series, preserving some ordering information in Transformers. Lastly, Mamba [8], a well-known sequential model from last year, uses a bidirectional Mamba block to extract inter-variate correlations and temporal dependencies.
5.3 Implementation Details
We adopted the default settings of the Time-Series-Library [36] to conduct a comprehensive comparison of baselines. We use the results of five random seeds as the average. We used 96 time steps as input and 336 time steps as ground truth. The code was implemented in PyTorch and executed on a V100 GPU with 32GB memory and 384GB RAM, provided by NCI Australia, an NCRIS-enabled capability supported by the Australian Government.
5.4 Results of Extremely Long Forecasting with gaps
We use Mean Squared Error (MSE) and Mean Absolute Error (MAE) metrics to evaluate performance, averaging results across different seeds. It is observed that nearly all results are poor, highlighting the significant challenge posed by domain shifts over time for extremely long forecasting with gaps. These baseline results also indicate that traditional SOTA rankings and methodologies are no longer effective. Notably, Autoformer, which performs the worst under conventional data and settings, shows the best performance in our setting. This is understandable because Autoformer’s core technology focuses on exploring the correlation within the data itself. This insight suggests that future efforts to tackle this problem should place greater emphasis on leveraging the intrinsic potential of the data.
Considering that our dataset provides an extensive temporal range for training, we can theoretically extend the input step length considerably. We extended the 96 steps input from Table 3 to a maximum of 1440 steps for testing. As shown in Table 4, the performance improves significantly with the increase in input step length, further demonstrating the substantial potential of our dataset for deep exploration.
Table 4: Results of ablation study with 4 different input step lengths
DLinear (Input Length) 96 192 336 720
Metrics MSE MAE MSE MAE MSE MAE MSE MAE
PEMS03_gap 1-year gap 96 1.500 0.933 1.455 0.912 1.462 0.906 1.392 0.887
192 1.542 0.945 1.147 0.910 1.457 0.906 1.445 0.906
336 1.531 0.935 1.447 0.907 1.462 0.906 1.439 0.904
5.5 Results of Hourly and Daily Forecasting
We believe that both the hourly and daily datasets are equally significant. Multi-scale, diverse datasets can provide the community with valuable references. We observe that the performance degrades progressively from the hourly to the daily to the gap datasets. Smaller time scales help reduce complexity and uncertainty, thereby improving prediction accuracy. Research has shown that clustering at different scales can enhance model performance [34]. Therefore, our aggregated version of the data will contribute new external features to the community.
Table 5: Comparison in hourly and daily datasets
Methods Mamba iTransformer DLinear Autoformer
Metrics MSE MAE MSE MAE MSE MAE MSE MAE
PEMS03_agg Hourly 96 0.144 0.222 0.530 0.535 0.159 0.222 0.241 0.346
192 0.173 0.237 0.215 0.289 0.153 0.208 0.235 0.340
336 0.158 0.220 0.519 0.527 0.167 0.216 0.260 0.362
Daily 96 0.754 0.503 0.606 0.419 0.602 0.426 0.771 0.537
192 0.968 0.604 0.781 0.500 0.794 0.509 0.897 0.577
336 1.210 0.706 0.967 0.579 0.984 0.584 1.058 0.630
PEMS04_agg Hourly 96 0.137 0.244 0.240 0.339 0.161 0.245 0.178 0.295
192 0.132 0.239 0.260 0.361 0.142 0.223 0.175 0.288
336 0.121 0.216 0.226 0.413 0.145 0.226 0.197 0.316
Daily 96 0.672 0.499 0.534 0.442 0.507 0.415 0.644 0.506
192 0.720 0.549 0.634 0.508 0.610 0.483 0.749 0.580
336 0.795 0.602 0.706 0.555 0.663 0.522 0.728 0.569
PEMS07_agg Hourly 96 0.212 0.302 0.375 0.425 0.203 0.259 0.307 0.390
192 0.201 0.288 0.297 0.368 0.182 0.231 0.313 0.391
336 0.191 0.245 0.126 0.156 0.190 0.241 0.291 0.364
Daily 96 1.719 0.736 1.426 0.613 1.414 0.606 1.703 0.762
192 2.005 0.842 1.772 0.730 1.756 0.720 1.903 0.804
336 2.290 0.949 2.078 0.819 2.051 0.813 2.171 0.884
PEMS08_agg Hourly 96 0.245 0.287 0.363 0.379 0.253 0.272 0.305 0.377
192 0.269 0.292 0.341 0.354 0.254 0.259 0.340 0.401
336 0.283 0.298 0.369 0.369 0.281 0.272 0.452 0.468
Daily 96 0.870 0.558 0.766 0.486 0.746 0.478 0.913 0.604
192 1.023 0.635 0.911 0.554 0.906 0.548 1.026 0.648
336 1.161 0.697 1.024 0.617 1.022 0.602 1.127 0.689
6 Prospects and Constraints
Prospects. Our dataset spans the longest time period among existing datasets, and it is not only the largest spatially, but also evolving in growth. It can continue to update in line with the updates from the PeMS system in the future. It is specifically designed for various complex scenarios, such as those already mentioned, including extremely long forecasting with long gaps, and hourly and daily predictions. Additionally, zero-shot forecasting designed for evolving growth scenarios will also be included in Appendix.
Constraints. The limitations of our dataset are also quite evident. Due to the sheer size of our dataset, it requires more computational resources. However, with the maturation of large language models and foundational models, we believe its large volume will become an advantage, contributing more diverse data to the spatio-temporal large model community.
Acknowledgments and Disclosure of Funding
We would like to acknowledge the support of Cisco’s National Industry Innovation Network (NIIN) Research Chair Program and the ARC Centre of Excellence for Automated Decision-Making and Society (CE200100005). We acknowledge the resources and services from the National Computational Infrastructure (NCI), which is supported by the Australian Government.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] See Section 1
    2. (b)
      Did you describe the limitations of your work? [Yes] See Section 6
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] We have discussed and identified no negative societal impact.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] See Section 4.1
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See Section 5.1 and Section 5.3
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] As discussed in Section 5.3, our experiments area averaged over five random seeds.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Section 5.3
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] We have cited the creators in Section 4.1
    2. (b)
      Did you mention the license of the assets? [Yes] See Section 4.4. Our dataset complies with the CC BY-NC 4.0 license, and the code adheres to the MIT license.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [No]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] Please refer to Section 4.1 for the detailed discussion.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [No] This dataset is distributed across the highway, collecting traffic speed, traffic flow, and traffic occupancy. It does not contain personally identifiable information or offensive content.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
Appendix A Appendix
A.1 Dataset
To support our setting of extremely long forecasting with gaps, we selected a subset installed in the early stages and have consistently collected new data up to the present, which is shown as follows:
Table 6: Overview of Gap, Hourly, and Daily Aggregated Data Based on Processed Raw Data
Datasets(Gap/Hour/Day) Time Period Nodes
PEMS03_gap&agg 03/2001 - 03/2024 151
PEMS04_gap&agg 06/2002 - 03/2024 822
PEMS05_gap&agg 03/2012 - 03/2024 103
PEMS06_gap&agg 12/2009 - 03/2024 130
PEMS07_gap&agg 06/2002 - 03/2024 3062
PEMS08_gap&agg 03/2001 - 03/2024 212
PEMS10_gap&agg 06/2007 - 03/2024 107
PEMS11_gap&agg 09/2002 - 03/2024 521
PEMS12_gap&agg 01/2002 - 03/2024 1543
A.2 Results
Here we provided the results of District 5,6,10,11,12, as shown in Table 7 and Table 8:
Table 7: Comparison in hourly and daily datasets in District 05,06,10,11,12. The symbol -- indicates that the result is an outlier.
Methods Mamba iTransformer DLinear Autoformer
Metrics MSE MAE MSE MAE MSE MAE MSE MAE
PEMS05_agg Hourly 96 0.121 0.205 0.226 0.324 0.148 0.236 0.155 0.272
192 0.118 0.197 0.214 0.312 0.132 0.216 0.159 0.268
336 0.115 0.194 0.220 0.318 0.134 0.217 0.167 0.274
Daily 96 0.655 0.511 0.654 0.510 0.607 0.477 0.650 0.522
192 0.798 0.598 0.775 0.576 0.698 0.535 0.643 0.512
336 0.907 0.614 0.787 0.582 0.745 0.556 0.764 0.578
PEMS06_agg Hourly 96 0.142 0.234 0.269 0.336 0.188 0.254 0.164 0.271
192 0.138 0.218 0.256 0.326 0.166 0.227 0.188 0.291
336 0.137 0.207 0.226 0.413 0.171 0.227 0.197 0.316
Daily 96 0.516 0.419 0.414 0.389 0.405 0.340 0.518 0.437
192 0.642 0.485 0.543 0.468 0.510 0.395 0.580 0.460
336 0.734 0.519 0.675 0.492 0.596 0.437 0.658 0.481
PEMS10_agg Hourly 96 0.213 0.256 0.391 0.413 0.272 0.309 0.260 0.346
192 0.205 0.250 0.387 0.412 0.246 0.277 0.380 0.417
336 0.211 0.255 0.394 0.413 0.258 0.281 0.329 0.386
Daily 96 1.161 0.671 0.926 0.513 0.951 0.567 1.079 0.647
192 1.459 0.784 1.414 0.730 1.228 0.681 1.429 0.786
336 1.715 0.855 1.478 0.768 1.451 0.751 1.552 0.817
PEMS11_agg Hourly 96 - 0.472 - 0.538 - 0.306 - 0.800
192 - 0.470 - 0.443 - 0.291 - 0.742
336 - 0.487 - 0.435 - 0.300 - 0.745
Daily 96 - - - - - - - -
192 - - - - - - - -
336 - - - - - - - -
PEMS12_agg Hourly 96 0.145 0.237 0.083 0.157 0.174 0.242 0.188 0.287
192 0.142 0.226 0.091 0.159 0.154 0.216 0.196 0.289
336 0.146 0.217 0.104 0.170 0.162 0.219 0.209 0.298
Daily 96 1.373 0.621 1.456 0.622 1.052 0.510 1.348 0.649
192 1.722 0.726 1.675 0.678 1.403 0.611 1.548 0.686
336 2.066 0.823 1.984 0.641 1.654 0.675 1.801 0.746
Table 8: Comparison in gap dataset in District 05,06,10,11,12. The bold text indicates the best.
Methods Mamba iTransformer DLinear Autoformer
Metrics MSE MAE MSE MAE MSE MAE MSE MAE
PEMS05_gap 1-year gap 96 2.079 1.209 1.945 1.164 1.291 0.916 1.065 0.796
192 2.132 1.256 1.984 1.185 1.750 1.099 1.063 0.809
336 2.377 1.340 2.067 1.234 1.894 1.144 1.135 0.827
1.5-year gap 96 1.852 1.122 1.879 1.078 1.683 1.054 1.060 0.785
192 1.929 1.182 1.593 1.032 1.633 1.045 0.912 0.712
336 2.370 1.313 2.214 1.071 0.794 1.184 1.345 0.882
2-year gap 96 1.868 1.106 1.580 0.969 1.602 1.018 0.828 0.672
192 2.219 1.274 1.481 0.958 1.589 1.027 1.018 0.772
336 2.695 1.212 2.207 1.201 1.922 1.139 1.186 0.839
PEMS06_gap 1-year gap 96 1.806 1.066 0.875 0.692 1.173 0.837 1.216 0.859
192 1.928 1.112 1.227 0.848 1.410 0.942 0.961 0.751
336 2.181 1.212 1.594 1.003 1.501 0.976 0.992 0.769
1.5-year gap 96 1.549 0.997 1.331 0.891 1.484 0.963 0.885 0.710
192 1.746 1.054 1.077 0.778 1.353 0.920 1.010 0.768
336 1.605 1.018 1.500 0.961 1.587 1.011 0.955 0.739
2-year gap 96 1.226 0.851 1.864 1.106 1.691 1.033 1.013 0.768
192 0.949 0.720 1.343 0.879 1.259 0.858 0.853 0.691
336 0.945 0.710 1.550 0.970 1.415 0.934 0.955 0.739
PEMS10_gap 1-year gap 96 2.310 1.233 0.882 0.698 1.466 0.963 1.208 0.865
192 2.878 1.396 1.606 0.993 1.834 1.106 1.152 0.860
336 2.584 1.327 1.413 0.825 1.887 1.121 1.130 0.830
1.5-year gap 96 2.181 1.206 1.812 1.052 1.765 1.057 1.368 0.913
192 2.525 1.318 1.645 0.986 1.791 1.069 1.151 0.828
336 2.488 1.310 2.051 1.134 1.976 1.139 1.111 0.809
2-year gap 96 1.165 0.822 2.772 1.383 1.977 1.119 1.188 0.845
192 1.082 0.792 1.974 1.084 1.490 0.932 0.971 0.751
336 1.021 0.764 2.157 1.168 1.619 0.989 1.096 0.809
PEMS11_gap 1-year gap 96 5.199 0.891 5.417 0.997 5.276 0.936 4.930 0.854
192 5.300 0.920 5.504 1.029 5.393 0.974 5.114 0.867
336 5.251 0.901 5.410 0.985 5.364 0.955 4.830 0.849
1.5-year gap 96 5.871 0.968 6.045 1.296 5.914 1.224 5.691 0.878
192 5.968 1.012 6.121 1.314 5.993 1.229 5.792 0.884
336 6.167 1.074 6.214 1.326 6.025 1.299 5.947 0.967
2-year gap 96 5.914 0.996 6.136 1.318 5.945 1.243 5.761 0.978
192 6.541 1.043 6.213 1.327 6.014 1.289 6.245 1.001
336 6.541 1.086 6.221 1.332 6.024 1.300 6.268 1.024
PEMS12_gap 1-year gap 96 1.751 1.025 1.624 1.002 1.611 1.005 1.060 0.789
192 1.726 1.024 1.424 0.929 1.537 0.972 1.150 0.840
336 1.751 1.029 1.672 1.015 1.683 1.017 0.889 0.719
1.5-year gap 96 1.554 0.967 1.479 0.921 1.468 0.910 0.954 0.875
192 1.401 0.898 1.314 0.867 1.301 0.862 0.943 0.869
336 1.417 0.906 1.322 0.869 1.298 0.859 0.921 0.846
2-year gap 96 0.956 0.704 0.876 0.671 0.872 0.664 0.846 0.659
192 0.846 0.656 0.813 0.653 0.806 0.649 0.785 0.628
336 0.814 0.628 0.789 0.631 0.776 0.624 0.754 0.617
Additionally, we provided the results of naive zero-shot forecasting, as shown in Table 9. The poor performance of this method indicates significant potential for improvement.
Table 9: Results of zero shot forecasting in PEMS03. In this study, we utilized the inputs from 151 existing nodes in PEMS03 to predict 52 new nodes, resulting in a test set comprising a total of 203 nodes.
Datasets 96 192 336
Metrics MSE MAE MSE MAE MSE MAE
PEMS03_gap 1-year gap 4.658 3.465 4.945 3.648 5.198 3.892
1.5-year gap 4.891 3.657 5.124 3.842 5.263 3.996
2-year gap 4.547 3.410 4.895 3.539 5.103 3.758

### v2

@@
-XXLTraffic: Expanding and Extremely Long Traffic Dataset for Ultra-Dynamic Forecasting Challenges
-Abstract
+XXLTraffic: Expanding and Extremely Long Traffic forecasting beyond test adaptation
+Abstract.
 Traffic forecasting is crucial for smart cities and intelligent transportation initiatives, where deep learning has made significant progress in modeling complex spatio-temporal patterns in recent years.
-However, current public datasets have limitations in reflecting the ultra-dynamic nature of real-world scenarios, characterized by continuously evolving infrastructures, varying temporal distributions, and temporal gaps due to sensor downtimes or changes in traffic patterns.
+However, current public datasets have limitations in reflecting the distribution shift nature of real-world scenarios, characterized by continuously evolving infrastructures, varying temporal distributions, and long temporal gaps due to sensor downtimes or changes in traffic patterns.
 These limitations inevitably restrict the practical applicability of existing traffic forecasting datasets.
-To bridge this gap, we present XXLTraffic, the largest available public traffic dataset with the longest timespan and increasing number of sensor nodes over the multiple years observed in the data, curated to support research in ultra-dynamic forecasting.
+To bridge this gap, we present XXLTraffic, largest available public traffic dataset with the longest timespan collected from Los Angeles, USA, and New South Wales, Australia, curated to support research in extremely long forecasting beyond test adaptation.
 Our benchmark includes both typical time-series forecasting settings with hourly and daily aggregated data and novel configurations that introduce gaps and down-sample the training size to better simulate practical constraints.
 We anticipate the new XXLTraffic will provide a fresh perspective for the time-series and traffic forecasting communities.
-It would also offer a robust platform for developing and evaluating models designed to tackle ultra-dynamic and extremely long forecasting problems.
+It would also offer a robust platform for developing and evaluating models designed to tackle the extremely long forecasting problems beyond test adaptation.
 Our dataset supplements existing spatio-temporal data resources and leads to new research directions in this domain.
-1 Introduction
-Traffic prediction plays a vital role in intelligent transportation systems by offering accurate and dependable forecasts of future traffic conditions that serve as critical references for route design, urban construction, and road management.
-The long-term spatio-temporal relationship is essential and irreplaceable for traffic forecasting, i.e., the long-term relationship can not only provide context that helps understand anomalies or irregularities in short-term data.
-but also help identify patterns and trends over different timescales, such as daily, weekly, monthly, and yearly cycles.
-In recent years, significant work has focused on short-term and long-term traffic flow prediction.
-Various deep learning techniques, such as GNNs, emphasize extracting spatial relationships in traffic networks [13], while Transformer variants and upgraded architectures and methods focus on capturing short-term and long-term temporal relationships [29].
-Existing work has achieved promising performance in both short-term and long-term predictions, but this is limited to relatively simple experimental setups.
-The inevitable tendency in intelligent transportation systems is to design models that incorporate more realistic circumstances rather than just simple experimental conditions.
-This requires the use of data sets that accurately reflect real-world scenarios.
-This serves as a strong incentive for us to create and develop XXLTraffic, i.e., why does XXLTraffic focus on ultra-dynamic forecasting?
-Why is it necessary to include extremely long-term data?
-We will introduce the commonalities and developments of existing datasets, as well as the difficulties in establishing XXLTraffic more specifically.
-Figure 1: Our dataset are evolving and longer than existing datasets.
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06copyright: acmlicensed
+1. Introduction
+Rapid global population growth and vehicle proliferation have intensified urban traffic congestion.
+As cities expand and personal transportation reliance grows, strain on road networks leads to longer commutes, higher fuel consumption, and increased emissions.
+Accurate traffic prediction is vital for intelligent transportation systems, informing strategies to mitigate congestion and enhance mobility through improved route planning and urban development.
+Effective forecasting requires capturing long-term spatio-temporal relationships in traffic data.
+Long-term analysis provides context for anomalies in short-term patterns and reveals trends influenced by population cycles, seasonal shifts, and yearly vehicle usage changes.
+These insights are crucial for developing robust models that adapt to evolving urban traffic dynamics due to demographic and vehicular changes.
+In recent years, significant work has focused on both short-term and long-term traffic flow prediction.
+Deep learning techniques, including Graph Neural Networks (GNNs), have been employed to extract spatial relationships within traffic networks (Jin et al., 2023), while Transformer-based architectures have been utilized to capture temporal dependencies over various timescales (Shao et al., 2023a).
+Although these methods have shown promising results, they often rely on datasets that do not fully encapsulate the complexities introduced by rapid population growth and the surging number of vehicles, thus limiting their applicability to real-world scenarios.
+There is an emerging need in intelligent transportation systems to design predictive models that extend beyond test adaptation, effectively generalizing to real-world conditions that evolve over time.
+It is important to note that our concept of ’beyond test adaptation’ differs from ’test time adaptation’ (Guo et al., 2024) as illustrated in Fig.
+1 that shows the distinctions between them.
+This shift requires models that can handle the multifaceted impacts of demographic changes and vehicle proliferation without relying solely on adaptation to specific test datasets.
+To achieve this, it is essential to utilize datasets that accurately represent these evolving conditions over extremely long periods, capturing the intricate patterns influenced by population and vehicular growth.
+Figure 1.
+Test-time adaptation in time-series forecasting involves training a single model to fit different test domains, horizons, or gaps.
+The figure above illustrates this using a gap example.
+In contrast, the figure below shows our ’beyond test adaptation’ where we train separate models for various gap settings.
+Motivated by this need, we introduce XXLTraffic, a dataset and framework that expands traffic forecasting beyond test adaptation.
+By incorporating extremely long-term data, XXLTraffic better reflects real-world scenarios where traffic patterns are continually affected not just by infrastructure changes like highway construction, but also by shifts in distribution due to factors like population growth and increasing vehicle numbers.
+We will discuss existing datasets and the specific challenges encountered in establishing XXLTraffic, highlighting how it advances the field by providing a more realistic and comprehensive dataset.
+This facilitates the development of models capable of adapting to the complexities of real-world traffic dynamics without the limitations of traditional test adaptation approaches.
+Figure 2.
+Our dataset is evolving and longer than existing datasets.
 Existing datasets are either limited by short temporal spans or insufficient spatial nodes.
 In contrast, our dataset features an evolving growth of spatial nodes and spans over 20 years.
-Recent Advances in Expanding Traffic Datasets.
+1.1. Recent Advances in Expanding Traffic Datasets
 Real-world traffic scenarios necessitate more complex prediction settings, involving extended temporal horizons or broader spatial coverage in experiments.
-In the temporal domain, new settings are typically proposed based on previously published work rather than introducing new datasets:  [30] and [11] expanded input and output lengths up to four times on existing datasets.From a spatial perspective, [4] published a dataset with nodes growing annually and provided an evolving network to support new node predictions.
-[33] proposed a continual learning framework with pattern expansion mechanisms based on [4].
-Additionally, [27] and [25] offered larger-scale spatial node datasets to support subsequent researchers.
+In the temporal domain, new settings are typically proposed based on previously published work rather than introducing new datasets:  (Shao et al., 2023b) and (Jia et al., 2024) expanded input and output lengths up to four times on existing datasets.From a spatial perspective, (Chen et al., 2021) published a dataset with nodes growing annually and provided an evolving network to support new node predictions.
+(Wang et al., 2023b) proposed a continual learning framework with pattern expansion mechanisms based on (Chen et al., 2021).
+Additionally, SCPT (Prabowo et al., 2024) and Large-ST (Liu et al., 2024b) offered larger-scale spatial node datasets to support subsequent researchers.
 Recent work has explored longer temporal step experimental settings and released traffic datasets spanning up to five years and thousands of nodes.
 However, in specific scenarios, such as future traffic prediction for highway planning, these data and experimental settings fall short.
-As shown in Figure 1, most existing datasets have limitations in temporal span, which inspired us to develop a dataset for expanding and extremely long traffic forecasting.
-Highway construction typically spans several years, requiring predictions several years in advance.
-These extremely long prediction scenarios necessitate traffic datasets with even longer temporal spans for support.
+As shown in Fig.
+2, most existing datasets have limitations in temporal span, which inspired us to develop a dataset for expanding and extremely long traffic forecasting.
+This need for traffic forecasting beyond test adaptation is crucial in various real-world scenarios.
+For instance, urban planning and infrastructure investment decisions rely heavily on accurate long-term traffic predictions to ensure that developments meet future transportation demands.
+Commercial real estate site selection and development also depend on knowing future traffic volumes years in advance to optimize location choices and investment strategies.
+Additionally, governments can formulate more effective environmental policies based on long-term traffic forecasts, such as implementing traffic restrictions or promoting electric vehicles to reduce emissions.
+These applications highlight the importance of developing predictive models capable of accurately forecasting traffic patterns over extended periods, facilitating strategic decision-making across multiple sectors.
 As the temporal span extends, urban infrastructure development and road construction can lead to shifts in traffic patterns, resulting in an evolving domain shift.
@@
 Additionally, the combination of these factors enables the extraction of more temporal patterns from extremely long sequences, allowing for the possibility of longer input sequences.
-Challenges.
+1.2. Challenges
 The ultra-dynamic challenge encompasses three key aspects: (1) Continuously evolving states of the underlying spatio-temporal infrastructures, characterized by an expanding number of nodes over the years.
 This continuous growth introduces complexity as the infrastructure adapts and expands.
-(2) Evolving temporal distributions over an extremely long observation period, which is crucial for extremely long forecasting.
+(2) Evolving temporal distributions over an extremely long observation period, which is crucial for extremely long forecasting beyond different noncontiguous train-test splits.
 This requires models to adapt to changes in patterns and trends over extensive temporal spans.
-(3) Ultra spatio-temporal dynamicity with temporal gaps.
-These gaps can occur due to various reasons such as sensor data becoming unavailable during road construction, maintenance activities, or shifts in data distributions.
-This dynamic nature necessitates robust modeling techniques to handle inconsistencies and interruptions in data.
-Contribution.
 We have constructed a traffic dataset with an exceptionally long temporal span and broader regional coverage, providing aggregated data and benchmarking, as well as a benchmarking setup considering extremely long prediction scenarios for future exploration:
@@
 This dataset supports both temporally scalable and spatially scalable challenges in traffic prediction.
-  • We present an experimental setup with temporal gaps for extremely long prediction with gaps scenarios and provide a benchmark of aggregated versions of hourly and daily datasets.
-  • We provide the exploration of input features through evolving temporal distributions over an extremely long observation period.
-Additionally, our datasets support zero-shot forecasting for new sensors.
-2 Preliminaries
+  • We present an experimental setup with temporal gaps for extremely long prediction beyong test adaptation and provide a benchmark of aggregated versions of hourly and daily datasets.
+  • We provide exploration of input features through evolving temporal distributions over an extremely long observation period.
+2. Preliminaries
 In this section, we will define traffic data and traffic prediction tasks.
@@
 Traffic data primarily consists of vehicle flow detection data collected by sensors distributed across various locations in the traffic network.
-It is generally represented by $X_{i}$in $\mathbb{R}^{N\times T\times C}$, where $T$ denotes the time steps, $N$ denotes the number of sensors, and $C$ denotes the number of features.
+It is generally represented by $X_{i}$in $\mathbb{R}^{N\times T\times C}$, where $T$ denotes the time steps, $N$ denotes the number of sensors and $C$ denotes the number of features.
 Definition 2. Short-term traffic prediction:
@@
 As shown in Equation 1, the input length $\alpha$ and output length $\beta$ are generally set to 12 steps.
-$[X_{t-(\alpha-1)},...,X_{t-1},X_{t}]\rightarrow[X_{t+1},X_{t+2},...,X_{t+\beta}],$ (1)
+(1) $[X_{t-(\alpha-1)},...,X_{t-1},X_{t}]\rightarrow[X_{t+1},X_{t+2},...,X_{t+\beta}],$
 Definition 3. Long-term multivariate prediction:
-This task mainly focuses on long-sequence time series prediction, which includes the traffic dataset. As shown Table 1, the sequence length can reach up to 2880 steps.
+This task mainly focuses on long-sequence time series prediction, which includes the traffic dataset. As shown in Table 1, the sequence length can reach up to 2880 steps.
 Definition 4. Extremely Long Prediction with Gaps:
-Based on Equation 1, the observation and prediction are not adjacent but are instead separated by a gap period $g$, as shown in the following formula.
-$[X_{t-(\alpha-1)},...,X_{t-1},X_{t}]\rightarrow[X_{t+g+1},X_{t+g+2},...,X_{t+g+\beta}],$ (2)
-3 Gaps and Comparison with Existing Traffic Datasets
-Figure 2: Our expanding and extremely long prediction addresses the existing limitations in both short-term and long-term predictions.
+Based on Equation 1, the observation and prediction are not adjacent but instead are separated by a gap period $g$, as shown in the following formula.
+(2) $[X_{t-(\alpha-1)},...,X_{t-1},X_{t}]\rightarrow[X_{t+g+1},X_{t+g+2},...,X_{t+g+\beta}],$
+3. Gaps and Comparison with Existing Traffic Datasets
+Figure 3. Our expanding and extremely long prediction addresses the existing limitations in both short-term and long-term predictions.
 As shown in Table 1, existing traffic prediction work can easily be divided into short-term and long-term settings.
-The short-term setting originated from the STGCN[41] work, while the long-term setting was first introduced by LSTNet [14] and subsequently established as a widely adopted experimental framework by Informer [45].
+The short-term setting originated from the STGCN(Yu et al., 2018) work, while the long-term setting was first introduced by LSTNet (Lai et al., 2018) and subsequently established as a widely adopted experimental framework by Informer (Zhou et al., 2021).
 In recent years, short-term prediction typically has a maximum step length of 12 steps, while long-term prediction reaches up to 720 steps.
-However, works such as Witran [11] and DAN [19] recognized the need for even longer step predictions in practical applications, extending the length to a maximum of four times the typical length.
+However, works such as Witran (Jia et al., 2024) and DAN (Li et al., 2024) recognized the need for even longer step predictions in practical applications, extending the length to a maximum of four times the typical length.
 Despite the differences in step lengths, their observed and predicted values are concatenated tightly together, as shown in Equation 1.
@@
 Typically, existing datasets lack the temporal coverage required to support gaps exceeding one year.At the same time, predicting several years in advance also implies the need to forecast traffic for sensors at new locations, taking into account the evolving nature of the road network.
-Even when such coverage is available, works like [33] and [4] utilize evolving datasets but do not provide sufficient data to train models for extended durations.
+Even when such coverage is available, works like (Wang et al., 2023b) and (Chen et al., 2021) utilize evolving datasets but do not provide sufficient data to train models for extended durations.
 To overcome these gaps, our Expanding and Extremely Long Traffic Dataset robustly supports these complex scenarios.
-Table 1: Summary of recent short-term traffic forecasting and long-term multivariate forecasting
+Table 1. Summary of recent short-term traffic forecasting and long-term multivariate forecasting
 Datasets Model Series Length
-Short-term STGCN [41] {3,6,9,12}
-DCRNN [20], GWN [40], BTF [3], DMSTGCN [10], {3,6,12}
-GTS [28], STGODE [7], PM-MemNet [16], STAEFormer [23]
-AGCRN [1], STSGCN [32], ,DSTAGNN [15] D2STGNN [31], {12}
-DyHSL [44],PDFormer [12], MultiSPANS [47], GMSDR [22]
-Long-term MTGNN [39],LSTNet [14] {3,6,12,24}
-ARU [6] {12,24,48,168,336}
-LogSparse_Trans [18] {24,48,72,96,120,144,168,192}
-AST [38] {8,24,168,336}
-SSDNet [21] {20,24,30,138}
-Informer [45], Autoformer [37], FEDformer [46], {24,48,96,192,336,720}
-Linear [17], Triformer [5], Pyraformer [24]
-DSformer [42],DeepTime [35],DLinear [43]
-Witran [11] {168, 336, 720, 1440, 2880}
-DAN [19] {288, 672, 1440}
-4 The XXLTraffic Datasets
-4.1 Data Collection
-We obtained the expanding and extremely long traffic sensor data from the California Department of Transportation (CalTrans) Performance Measurement System [^1] https://pems.dot.ca.gov/ (PeMS) [2].
+Short-term STGCN (Yu et al., 2018) {3,6,9,12}
+DCRNN (Li et al., 2018), GWN (Wu et al., 2019), BTF (Chen and Sun, 2021), {3,6,12}
+DMSTGCN (Han et al., 2021), GTS (Shang et al., 2021), STGODE (Fang et al., 2021),
+PM-MemNet (Lee et al., 2021), STAEFormer (Liu et al., 2023)
+AGCRN (Bai et al., 2020), STSGCN (Song et al., 2020), ,DSTAGNN (Lan et al., 2022), {12}
+D2STGNN (Shao et al., 2022), DyHSL (Zhao et al., 2023),PDFormer (Jiang et al., 2023),
+MultiSPANS (Zou et al., 2024), GMSDR (Liu et al., 2022)
+Long-term MTGNN (Wu et al., 2020a),LSTNet (Lai et al., 2018) {3,6,12,24}
+ARU (Deshpande and Sarawagi, 2019) {12,24,48,168,336}
+LogSparse_Trans (Li et al., 2019) {24,48,72,96,120,144,168,192}
+AST (Wu et al., 2020b) {8,24,168,336}
+SSDNet (Lin et al., 2021) {20,24,30,138}
+Informer (Zhou et al., 2021), Autoformer (Wu et al., 2021), FEDformer (Zhou et al., 2022), {24,48,96,192,336,720}
+Linear (Li et al., 2022), Triformer (Cirstea et al., 2022), Pyraformer (Liu et al., 2021)
+DSformer (Yu et al., 2023),DeepTime (Woo et al., 2023),DLinear (Zeng et al., 2023)
+Witran (Jia et al., 2024) {168, 336, 720, 1440, 2880}
+DAN (Li et al., 2024) {288, 672, 1440}
+4. The XXLTraffic Datasets
+4.1. Data Collection
+We obtained the expanding and extremely long traffic sensor data from the California Department of Transportation (CalTrans) Performance Measurement System [^1] https://pems.dot.ca.gov/ (PeMS) (Chen et al., 2001) and Transport for NSW [^2] https://maps.transport.nsw.gov.au/egeomaps/traffic-volumes/index.html.
 PeMS is an online platform that collects traffic data from 19,561 sensors distributed across California state highways.
@@
 We have stored the complete raw data files in an open-source repository for quick access, which will be released after the publication.
-Table 2: Comparison between our XXLTraffic dataset and the existing traffic datasets.
+The tfNSW is an open-source data platform provided by Transport for NSW, featuring traffic flow data collected from sensors distributed along major roads throughout the state of New South Wales of Australia.
+The data is available at a minimum granularity of one hour.
+Table 2. Comparison between our XXLTraffic dataset and the existing traffic datasets.
 Reference Dataset Samples Nodes Time Interval Time Span Time Period
-DCRNN [20] METR-LA 34,272 207 5 mins 4 months 03/2012 - 06/2012
+DCRNN METR-LA 34,272 207 5 mins 4 months 03/2012 - 06/2012
 PEMS-BAY 52,116 325 5 mins 6 months 01/2017 - 05/2017
-LSTNet [14] Traffic 17,544 862 1 hour 2 years 01/2015 - 12/2016
-STGCN [41] PEMSD7(M) 12,672 228 5 mins 2 months 05/2012 - 06/2012
+LSTNet Traffic 17,544 862 1 hour 2 years 01/2015 - 12/2016
+STGCN PEMSD7(M) 12,672 228 5 mins 2 months 05/2012 - 06/2012
 PEMSD7(L) 12,672 1026 5 mins 2 months 05/2012 - 06/2012
-ASTGCN [9] PEMSD4-I 17,002 228 5 mins 2 months 01/2018 - 02/2018
+ASTGCN PEMSD4-I 17,002 228 5 mins 2 months 01/2018 - 02/2018
 PEMSD8-I 17,856 1,979 5 mins 2 months 07/2016 - 08/2016
-STSGCN [32] PEMS03 26,208 358 5 mins 11 months 01/2018 - 11/2018
+STSGCN PEMS03 26,208 358 5 mins 11 months 01/2018 - 11/2018
 PEMS04 16,992 307 5 mins 2 months 01/2018 - 02/2018
@@
 PEMS08 17,856 170 5 mins 2 months 07/2016 - 08/2016
-Large-ST [25] CA 525,888 8,600 5 mins 5 years 01/2017 - 12/2021
+Large-ST CA 525,888 8,600 5 mins 5 years 01/2017 - 12/2021
 GLA 525,888 3,834 5 mins 5 years 01/2017 - 12/2021
@@
 Full_PEMS12 2,331,360 2,587 5 mins 22.16 years 01/2002 - 03/2024
+tfNSW 100,056 27 60 mins 11.42 years 01/2013 - 05/2024
 As illustrated in Table 2, our collected dataset significantly exceeds existing datasets in terms of both temporal coverage and the number of spatial nodes.
-The dataset sample will be available on: https://github.com/cruiseresearchgroup/XXLTraffic, which includes the raw data, sensor metadata (containing sensor IDs, geographical coordinates, associated road information, etc.), the data processing pipeline code, and the processed datasets.
-4.2 Data Preprocessing
+The dataset sample will be available on: https://anonymous.4open.science/r/XXLTraffic-F281, which includes the raw data, sensor metadata (containing sensor IDs, geographical coordinates, associated road information, etc.), the data processing pipeline code, and the processed datasets.
+4.2. Data Preprocessing
 Based on the 23 years of raw data we collected, we conducted rigorous data filtering and aggregation.
@@
 To support our setting of extremely long forecasting with gaps, we selected a subset of sensors that were installed in the early stages and have consistently collected new data up to the present(named gap dataset), which is shown in the Appendix.
-This extensive gap dataset effectively underpins the extremely long forecasting with gaps demonstrated in Figure 2.
+This extensive gap dataset effectively underpins the extremely long forecasting with gaps demonstrated in Fig.
+3.
 Utilizing the gap dataset, we performed both hourly and daily aggregations, which will be employed for gap-free long-term forecasting benchmarking.
 We will provide standard long-term forecasting benchmarks for both the hourly and daily datasets.
-4.3 Data Overview
-The XXLTraffic dataset is distributed across highways in the state of California, as illustrated in the Figure 3(a).
-The nine colors represent nine districts.
-From Figures 3(b),  3(c), and 3(d), we can clearly observe the evolutionary growth of the sensors.
-The sensors are extensively distributed across both urban and suburban areas, offering diverse modalities.
-Additionally, the sensors are densely interconnected, enabling the formation of a high-quality traffic graph dataset.
 (b) District 08 in 2005
@@
 (d) District 08 in 2024
-Figure 3: XXLTraffic dataset overview and its evolving development.
+Figure 4.
+XXLTraffic dataset overview and its evolving development.
 This figure provides a global overview and two local overviews, showcasing the diversity of sensor distribution.
 The lower parts highlights a selected region to illustrate the growth and changes in traffic sensors over time.
+4.3. Data Overview
+The XXLTraffic dataset is distributed across highways in the state of California, as illustrated in the Fig.
+4(a).
+The nine colors represent nine districts.
+From Figures.
+4(b),  4(c), and 4(d), we can clearly observe the evolutionary growth of the sensors.
+The sensors are extensively distributed across both urban and suburban areas, offering diverse modalities.
+Additionally, the sensors are densely interconnected, enabling the formation of a high-quality traffic graph dataset.
 It is evident that sensors at the same location may collect completely different distributions over the course of urban evolution.
-As shown in Figure 4, some sensors have maintained the same distribution from 2005 to 2024, while others have experienced significant changes in distribution.
+As shown in Fig.
+5, some sensors have maintained the same distribution from 2005 to 2024, while others have experienced significant changes in distribution.
 The temporal changes causing domain shifts present a significant challenge for our extremely long forecasting.
-Figure 4: Sensor traffic status distribution of District 8 from 2005 to 2024.
+(a) distribution of District 8 in PeMS.
+(b) distribution of tfNSW.
+Figure 5.
+Sensor traffic status distribution of District 8 in PeMS from 2005 to 2024 in 5(a) and from 2016 to 2022 in NSW in 5(b).
 While some sensors exhibit minimal changes, others show significant distribution differences, regardless of whether they are in low-traffic or high-traffic areas.
 This presents substantial challenges for extremely long forecasting with long gaps.
-4.4 XXLTraffic Licence
+4.4. XXLTraffic Licence
 The XXLTraffic dataset is licensed under CC BY-NC 4.0 International: https://creativecommons.org/licenses/by-nc/4.0.
@@
 Please check the official repositories for the licenses of any specific baseline methods used in our codebase.
-5 Experiments
+5. Experiments
 We conducted experiments for both extremely long forecasting with gaps using gap dataset and conventional long-term forecasting using hourly dataset and daily dataset.
-Additionally, referring to the definition in Figure 2, we set the gap parameter $g$ as 1 year, 1.5 years, and 2 years for the gap dataset, as illustrated by Figure 5.
-Figure 5: Problem definition. The yellow boxes represent typical predictions, the gray boxes denote gap periods between observation and prediction, and the blue boxes indicate extended predictions.
-5.1 Datasets
+Additionally, referring to the definition in Fig.
+3, we set the gap parameter $g$ as 1 year, 1.5 years, and 2 years for the gap dataset, as illustrated by Fig.
+6.
+Figure 6. Problem definition. The yellow boxes represent typical predictions, the gray boxes denote gap periods between observation and prediction, and the blue boxes indicate extended predictions.
+5.1. Datasets
 We conducted experiments on all proposed sub-datasets.
-To maintain consistency with previous state-of-the-art benchmarks, we selected districts 03, 04, 07, and 08 (widely recognized as PEMS03/04/07/08) for our experiments, using the gap, hourly, and daily datasets from these districts.
+To maintain consistency with previous state-of-the-art benchmarks, we selected districts 03, 04, and 08 (widely recognized as PEMS03/04/08) for the experiments using the gap, and districts 03, 04,07 and 08 for hourly, and daily experiments.
 Results for other datasets are presented in Appendix.
@@
 For the gap dataset, due to the extensive span of up to 20 years resulting in a large sample size, we fixed a seed during data preprocessing to select 10% of the dataset for training and testing to quickly demonstrate our results.
-In our experiments, due to GPU memory constraints, we pruned PEMS07_ours and PEMS12_ours by removing nodes with a missing data ratio greater than 10%, reducing the node number to 1,613 and 867.
 The details of the datasets used in our benchmarking is in Appendix A.1.
-Table 3: Comparison in gap dataset. The bold text indicates the best.
-Methods Mamba iTransformer DLinear Autoformer
-Metrics MSE MAE MSE MAE MSE MAE MSE MAE
-PEMS03_gap 1-year gap 96 1.457 0.913 1.636 0.989 1.500 0.933 1.301 0.906
-192 1.472 0.922 1.597 0.970 1.542 0.945 1.266 0.867
-336 1.434 0.913 1.512 0.935 1.531 0.935 1.137 0.807
-1.5-year gap 96 1.485 0.942 1.879 1.078 1.653 0.976 1.467 0.945
-192 1.442 0.934 1.753 1.017 1.642 0.968 1.460 0.923
-336 1.446 0.938 1.662 0.987 1.632 0.968 1.345 0.882
-2-year gap 96 1.359 0.833 1.844 1.048 1.568 0.954 1.328 0.894
-192 1.216 0.772 1.729 1.008 1.473 0.896 1.235 0.837
-336 1.294 0.801 1.614 0.960 1.407 0.870 1.220 0.838
-PEMS04_gap 1-year gap 96 1.325 0.914 1.644 1.037 1.396 0.955 0.819 0.717
-192 1.438 0.642 1.587 1.018 1.440 0.965 0.941 0.767
-336 1.424 0.935 1.447 0.960 1.421 0.954 0.853 0.730
-1.5-year gap 96 1.204 0.890 2.014 1.193 1.488 0.995 0.981 0.779
-192 0.982 0.787 1.649 1.038 1.301 0.913 0.867 0.735
-336 0.961 0.792 1.352 0.929 1.298 0.908 0.762 0.679
-2-year gap 96 1.220 0.893 1.652 1.074 1.446 0.977 0.909 0.755
-192 1.004 0.807 1.189 0.870 1.268 0.893 0.909 0.747
-336 1.198 0.864 1.584 1.016 1.269 0.898 0.898 0.736
-PEMS07_gap 1-year gap 96 1.719 1.006 1.756 1.024 1.680 1.024 1.215 0.867
-192 1.637 0.996 1.816 1.061 1.776 1.043 1.202 0.852
-336 1.637 0.991 1.784 1.028 1.774 1.034 1.098 0.774
-1.5-year gap 96 1.571 0.972 1.705 1.016 1.585 0.987 1.184 0.844
-192 1.578 0.982 1.657 1.004 1.624 0.996 1.088 0.800
-336 1.818 1.127 1.712 1.028 1.668 1.036 0.991 0.729
-2-year gap 96 5.411 1.319 2.055 1.149 1.746 1.048 1.099 0.802
-192 12.620 1.370 2.006 1.128 1.705 1.002 1.284 0.875
-336 9.614 1.378 1.812 1.057 1.605 0.968 0.972 0.734
-PEMS08_gap 1-year gap 96 5.411 1.319 1.514 0.950 1.771 1.058 1.153 0.843
-192 12.620 1.370 1.499 0.900 1.762 1.059 1.202 0.845
-336 9.614 1.378 1.654 0.984 1.961 1.131 1.184 0.849
-1.5-year gap 96 4.453 1.061 2.286 1.196 1.978 1.072 1.428 0.901
-192 9.413 1.046 1.713 0.971 1.606 0.928 1.314 0.833
-336 10.457 1.063 1.890 1.488 1.736 0.985 1.320 0.836
-2-year gap 96 5.117 0.953 2.400 1.232 1.969 1.069 1.474 0.895
-192 12.769 0.933 1.974 1.070 1.670 0.942 1.505 0.883
-336 13.382 0.946 1.936 1.056 1.628 0.926 1.464 0.874
-5.2 Baselines
+Table 3. Comparison in gap dataset. The bold text indicates the best.
+Gap Data Gap Metric Horizon Mamba iTrans DLinear Autofor Infor FEDFor MICN PatchTST STGCN ASTGCN GWN AGCRN PDFor
+PEMS03 1-year MSE 96 1.457 1.636 1.500 1.301 0.673 0.934 0.514 0.803 0.556 0.765 0.676 0.596 0.621
+192 1.472 1.597 1.542 1.266 0.739 0.960 0.528 0.852 0.562 0.764 0.581 0.565 0.545
+336 1.434 1.512 1.531 1.137 0.699 0.849 0.493 0.825 0.561 0.717 0.582 0.580 0.574
+MAE 96 0.913 0.989 0.933 0.906 0.552 0.697 0.515 0.647 0.536 0.618 0.574 0.562 0.576
+192 0.922 0.970 0.945 0.867 0.581 0.715 0.513 0.664 0.539 0.626 0.546 0.543 0.536
+336 0.913 0.935 0.935 0.807 0.560 0.640 0.491 0.636 0.538 0.598 0.543 0.552 0.548
+1.5-year MSE 96 1.485 1.879 1.653 1.467 1.138 1.190 0.839 1.245 1.256 1.441 1.250 1.168 1.256
+192 1.442 1.753 1.642 1.266 1.276 1.374 0.844 1.339 1.187 1.395 1.194 1.122 1.117
+336 1.446 1.662 1.632 1.137 1.340 1.259 0.755 1.359 1.167 1.273 1.015 1.176 1.182
+MAE 96 0.942 1.078 0.976 0.945 0.763 0.795 0.686 0.863 0.884 0.959 0.875 0.866 0.899
+192 0.934 1.017 0.968 0.867 0.827 0.715 0.681 0.906 0.872 0.947 0.856 0.843 0.842
+336 0.938 0.987 0.968 0.807 0.845 0.640 0.632 0.903 0.853 0.877 0.832 0.862 0.871
+2-year MSE 96 1.359 1.844 1.568 1.328 1.642 1.205 1.359 1.817 2.059 2.236 1.987 1.950 2.068
+192 1.216 1.729 1.473 1.235 1.955 1.489 1.308 1.899 2.055 2.180 1.877 1.746 1.897
+336 1.294 1.614 1.407 1.220 1.982 1.615 0.966 1.656 2.022 1.889 1.847 1.930 1.982
+MAE 96 0.833 1.048 0.954 0.894 0.956 0.816 0.911 1.104 1.176 1.234 1.148 1.150 1.189
+192 0.772 1.008 0.896 0.837 1.067 0.933 0.866 1.119 1.184 1.213 1.117 1.070 1.131
+336 0.801 0.960 0.870 0.838 1.086 0.978 0.700 1.016 1.169 1.117 1.107 1.140 1.161
+PEMS04 1-year MSE 96 1.325 1.644 1.396 0.819 0.624 0.712 0.721 1.309 1.233 1.311 1.310 1.290 1.101
+192 1.438 1.587 1.440 0.941 0.694 0.679 0.611 1.358 1.398 1.443 1.255 1.318 1.078
+336 1.424 1.447 1.421 0.853 0.668 0.584 0.526 1.338 1.505 1.325 1.292 1.302 1.085
+MAE 96 0.914 1.037 0.955 0.717 0.593 0.650 0.632 0.917 0.904 0.929 0.870 0.865 0.851
+192 0.642 1.018 0.965 0.767 0.634 0.638 0.586 0.932 1.025 1.023 0.879 0.907 0.816
+336 0.935 0.960 0.954 0.730 0.615 0.592 0.534 0.916 1.104 0.938 0.865 0.899 0.821
+1.5-year MSE 96 1.204 2.014 1.488 0.981 0.618 0.664 0.638 1.171 1.369 1.350 1.684 1.222 1.097
+192 0.982 1.649 1.301 0.867 0.622 0.679 0.570 1.109 1.642 1.549 1.501 1.346 1.214
+336 0.961 1.352 1.298 0.762 0.646 0.488 0.482 1.158 1.368 1.199 1.584 0.231 1.130
+MAE 96 0.890 1.193 0.995 0.779 0.592 0.632 0.581 0.870 1.082 0.937 1.063 0.961 0.849
+192 0.787 1.038 0.913 0.735 0.588 0.638 0.547 0.846 1.035 1.075 0.988 0.989 0.903
+336 0.792 0.929 0.908 0.679 0.598 0.527 0.494 0.850 0.862 0.832 0.997 0.963 0.854
+2-year MSE 96 1.220 1.652 1.446 0.909 0.650 0.685 0.666 1.284 1.653 1.247 1.669 1.236 1.099
+192 1.004 1.189 1.268 0.909 0.639 0.621 0.596 1.151 1.545 1.356 1.554 1.269 1.159
+336 1.198 1.584 1.269 0.898 0.717 0.521 0.481 1.205 1.556 1.174 1.128 1.314 1.032
+MAE 96 0.893 1.074 0.977 0.755 0.604 0.644 0.609 0.913 1.074 0.894 1.057 0.901 0.861
+192 0.807 0.870 0.893 0.747 0.601 0.607 0.570 0.867 1.023 0.948 1.041 0.915 0.891
+336 0.864 1.016 0.898 0.736 0.640 0.552 0.502 0.881 1.011 0.842 0.869 0.947 0.816
+PEMS08 1-year MSE 96 5.411 1.514 1.771 1.153 - 1.636 1.556 1.926 2.531 3.119 2.185 2.158 1.921
+192 12.620 1.499 1.762 1.202 - 1.401 1.408 1.887 2.560 2.405 2.223 2.144 2.248
+336 9.614 1.654 1.961 1.184 - 1.870 1.093 1.962 2.482 2.086 2.228 2.036 1.994
+MAE 96 1.319 0.950 1.058 0.843 - 1.384 0.979 1.104 1.325 1.404 1.280 1.220 1.148
+192 1.370 0.900 1.059 0.845 - 0.877 0.918 1.106 1.337 1.276 1.311 1.213 1.260
+336 1.378 0.984 1.131 0.849 - 1.078 0.761 1.133 1.317 1.181 1.312 1.185 1.179
+1.5-year MSE 96 4.453 2.286 1.978 1.428 - 1.389 1.034 1.362 1.772 2.370 1.309 1.166 1.664
+192 9.413 1.713 1.606 1.314 - 1.179 1.039 1.666 1.522 1.730 1.074 1.144 1.518
+336 10.457 1.890 1.736 1.320 - 1.197 0.868 1.272 1.495 1.444 1.533 1.078 1.049
+MAE 96 1.061 1.196 1.072 0.901 - 0.871 0.758 0.867 1.084 1.192 0.896 0.808 1.055
+192 1.046 0.971 0.928 0.833 - 0.768 0.755 1.001 0.979 1.043 0.805 0.810 1.012
+336 1.063 1.488 0.985 0.836 - 0.782 0.680 0.853 0.966 0.948 0.999 0.784 0.801
+2-year MSE 96 5.117 2.400 1.969 1.474 - 1.494 1.030 1.393 1.336 1.962 1.789 1.393 1.071
+192 12.769 1.974 1.670 1.505 - 1.292 1.044 1.229 1.218 1.292 1.007 1.229 1.171
+336 13.382 1.936 1.628 1.464 - 1.253 0.904 1.246 1.181 1.227 1.181 1.246 1.003
+MAE 96 0.953 1.232 1.069 0.895 - 0.885 0.740 0.856 0.868 1.069 1.017 0.856 0.792
+192 0.933 1.070 0.942 0.883 - 0.795 0.741 0.813 0.819 0.795 0.761 0.813 0.826
+336 0.946 1.056 0.926 0.874 - 0.772 0.685 0.814 0.806 0.842 0.813 0.814 0.768
+tfNSW 1-year MSE 96 1.480 1.236 1.166 1.320 1.354 1.377 1.219 0.946 1.176 1.404 1.451 1.283 1.070
+192 1.543 1.088 1.184 1.599 1.229 1.312 1.262 1.026 1.495 1.547 1.580 1.267 1.019
+336 1.459 1.099 1.196 1.477 1.242 1.219 1.366 0.949 1.517 1.617 1.534 1.048 0.822
+MAE 96 0.845 0.822 0.748 0.879 0.777 0.895 0.778 0.778 0.776 0.874 0.871 0.824 0.756
+192 0.867 0.733 0.758 0.961 0.757 0.890 0.799 0.821 0.880 0.889 0.904 0.828 0.753
+336 0.847 0.739 0.763 0.934 0.776 0.838 0.833 0.774 0.901 0.910 0.892 0.786 0.710
+1.5-year MSE 96 1.745 1.355 1.372 1.658 1.615 1.440 1.204 1.025 1.220 1.286 1.598 1.249 1.153
+192 1.784 1.260 1.367 1.666 1.484 1.452 1.251 1.058 1.624 1.615 1.620 1.228 1.095
+336 1.628 1.284 1.385 1.515 1.500 1.346 1.203 0.995 1.621 1.621 1.220 1.163 1.049
+MAE 96 0.944 0.877 0.846 1.016 0.879 0.919 0.778 0.807 0.785 0.813 0.940 0.809 0.790
+192 0.968 0.818 0.846 0.997 0.854 0.932 0.789 0.823 0.934 0.902 0.950 0.808 0.792
+336 0.939 0.835 0.854 0.945 0.883 0.871 0.770 0.788 0.945 0.900 0.814 0.811 0.775
+2-year MSE 96 1.364 1.019 1.008 1.319 1.101 1.318 1.033 0.848 1.201 1.331 1.414 1.124 0.969
+192 1.195 0.884 1.024 1.319 1.054 1.222 1.094 0.909 1.287 1.061 1.109 1.058 0.941
+336 1.588 0.907 1.166 1.233 0.787 1.056 0.959 0.865 1.260 0.996 1.011 0.960 0.917
+MAE 96 0.857 0.782 0.742 0.907 0.713 0.879 0.745 0.727 0.785 0.823 0.862 0.749 0.702
+192 0.783 0.701 0.749 0.891 0.721 0.855 0.737 0.757 0.831 0.726 0.755 0.722 0.737
+336 0.903 0.730 0.748 0.849 0.639 0.766 0.691 0.714 0.810 0.693 0.695 0.749 0.715
+5.2. Baselines
 In our comparison experiments, we adopted four popular baselines, including MLP, Transformer, and Mamba architectures.
-Autoformer [37], an earlier state-of-the-art model, leverages a decomposition architecture and auto-correlation mechanism to enhance efficiency and accuracy in long-term time series forecasting, outperforming traditional Transformer models.
-iTransformer [26] is the latest and most effective Transformer-based model, utilizing attention and feed-forward networks on inverted dimensions, embedding time points into variate tokens.
-DLinear [43] challenges the effectiveness of Transformer models by proposing a simple one-layer linear model that captures temporal relations in an ordered set of continuous points.
+Informer (Zhou et al., 2021) introduces an efficient transformer for long sequence time-series forecasting using ProbSparse self-attention and self-attention distilling, enabling encoder-decoder architectures to handle long sequences effectively.
+MICN (Wang et al., 2023a) proposes a multi-scale context network that models both local and global contexts for long-term time series forecasting, capturing patterns across different temporal scales to enhance performance.
+FEDformer (Zhou et al., 2022) introduces a frequency-enhanced decomposed transformer that models time series in both time and frequency domains, improving long-term forecasting by effectively capturing temporal patterns.
+PatchTST (Nie et al., [n.
+d.]) applies transformers to time series by treating them as sequences of patches, enabling effective long-term forecasting through self-attention over patch representations to capture temporal dependencies.Autoformer (Wu et al., 2021), an earlier state-of-the-art model, leverages a decomposition architecture and auto-correlation mechanism to enhance efficiency and accuracy in long-term time series forecasting, outperforming traditional Transformer models.
+iTransformer (Liu et al., 2024a) is the latest and most effective Transformer-based model, utilizing attention and feed-forward networks on inverted dimensions, embedding time points into variate tokens.
+DLinear (Zeng et al., 2023) challenges the effectiveness of Transformer models by proposing a simple one-layer linear model that captures temporal relations in an ordered set of continuous points.
 It employs positional encoding and uses tokens to embed sub-series, preserving some ordering information in Transformers.
-Lastly, Mamba [8], a well-known sequential model from last year, uses a bidirectional Mamba block to extract inter-variate correlations and temporal dependencies.
-5.3 Implementation Details
-We adopted the default settings of the Time-Series-Library [36] to conduct a comprehensive comparison of baselines.
+Lastly, Mamba (Gu and Dao, 2023), a well-known sequential model from last year, uses a bidirectional Mamba block to extract inter-variate correlations and temporal dependencies.
+Additionally, we have selected five SOTA baselines  (Yu et al., 2018; Guo et al., 2019; Wu et al., 2019; Bai et al., 2020; Jiang et al., 2023) from traffic forecasting domain.
+5.3. Implementation Details
+We adopted the default settings of the Time-Series-Library (Wu et al., 2022) to conduct a comprehensive comparison of baselines.
 We use the results of five random seeds as the average.
@@
 The code was implemented in PyTorch and executed on a V100 GPU with 32GB memory and 384GB RAM, provided by NCI Australia, an NCRIS-enabled capability supported by the Australian Government.
-5.4 Results of Extremely Long Forecasting with gaps
+5.4. Results of Extremely Long Forecasting with gaps
 We use Mean Squared Error (MSE) and Mean Absolute Error (MAE) metrics to evaluate performance, averaging results across different seeds.
 It is observed that nearly all results are poor, highlighting the significant challenge posed by domain shifts over time for extremely long forecasting with gaps.
-These baseline results also indicate that traditional SOTA rankings and methodologies are no longer effective.
-Notably, Autoformer, which performs the worst under conventional data and settings, shows the best performance in our setting.
-This is understandable because Autoformer’s core technology focuses on exploring the correlation within the data itself.
+These baseline results also indicate that traditional the-state-of-the-art (SOTA) rankings and methodologies are no longer effective.
+Notably, MICN, which performs the worst under conventional data and settings, shows the best performance in our setting.
+This is understandable because MICN’s core technology focuses on exploring the correlation within the data itself.
 This insight suggests that future efforts to tackle this problem should place greater emphasis on leveraging the intrinsic potential of the data.
+Additionally, we have compiled the training time for one epoch of all the baselines on the largest dataset, PEMS04_gap, and the smallest dataset, tfNSW, as shown in Table 9.
 Considering that our dataset provides an extensive temporal range for training, we can theoretically extend the input step length considerably.
@@
 As shown in Table 4, the performance improves significantly with the increase in input step length, further demonstrating the substantial potential of our dataset for deep exploration.
-Table 4: Results of ablation study with 4 different input step lengths
+Table 4. Results of ablation study with 4 different input step lengths
 DLinear (Input Length) 96 192 336 720
@@
 336 1.531 0.935 1.447 0.907 1.462 0.906 1.439 0.904
-5.5 Results of Hourly and Daily Forecasting
+5.5. Results of Hourly and Daily Forecasting
 We believe that both the hourly and daily datasets are equally significant.
@@
 We observe that the performance degrades progressively from the hourly to the daily to the gap datasets.
-Smaller time scales help reduce complexity and uncertainty, thereby improving prediction accuracy.
-Research has shown that clustering at different scales can enhance model performance [34].
+Smaller time scales help reduce complexity and uncertainty, thereby improving the accuracy of prediction.
+Research has shown that clustering at different scales can improve model performance (Wang et al., 2024).
 Therefore, our aggregated version of the data will contribute new external features to the community.
-Table 5: Comparison in hourly and daily datasets
+Table 5. Comparison in hourly and daily datasets
 Methods Mamba iTransformer DLinear Autoformer
@@
 336 1.161 0.697 1.024 0.617 1.022 0.602 1.127 0.689
-6 Prospects and Constraints
+6. Prospects and Constraints
 Prospects.
 Our dataset spans the longest time period among existing datasets, and it is not only the largest spatially, but also evolving in growth.
-It can continue to update in line with the updates from the PeMS system in the future.
+It can continue to update according to the updates from the PeMS system in the future.
 It is specifically designed for various complex scenarios, such as those already mentioned, including extremely long forecasting with long gaps, and hourly and daily predictions.
-Additionally, zero-shot forecasting designed for evolving growth scenarios will also be included in Appendix.
 Constraints.
@@
 However, with the maturation of large language models and foundational models, we believe its large volume will become an advantage, contributing more diverse data to the spatio-temporal large model community.
-Acknowledgments and Disclosure of Funding
-We would like to acknowledge the support of Cisco’s National Industry Innovation Network (NIIN) Research Chair Program and the ARC Centre of Excellence for Automated Decision-Making and Society (CE200100005).
-We acknowledge the resources and services from the National Computational Infrastructure (NCI), which is supported by the Australian Government.
-Checklist
-  1. 1.
-    For all authors…
-    1. (a)
-      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] See Section 1
-    2. (b)
-      Did you describe the limitations of your work? [Yes] See Section 6
-    3. (c)
-      Did you discuss any potential negative societal impacts of your work? [Yes] We have discussed and identified no negative societal impact.
-    4. (d)
-      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
-  2. 2.
-    If you are including theoretical results…
-    1. (a)
-      Did you state the full set of assumptions of all theoretical results? [N/A]
-    2. (b)
-      Did you include complete proofs of all theoretical results? [N/A]
-  3. 3.
-    If you ran experiments (e.g. for benchmarks)…
-    1. (a)
-      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] See Section 4.1
-    2. (b)
-      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See Section 5.1 and Section 5.3
-    3. (c)
-      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)?
-[Yes] As discussed in Section 5.3, our experiments area averaged over five random seeds.
-    4. (d)
-      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Section 5.3
-  4. 4.
-    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
-    1. (a)
-      If your work uses existing assets, did you cite the creators? [Yes] We have cited the creators in Section 4.1
-    2. (b)
-      Did you mention the license of the assets? [Yes] See Section 4.4. Our dataset complies with the CC BY-NC 4.0 license, and the code adheres to the MIT license.
-    3. (c)
-      Did you include any new assets either in the supplemental material or as a URL? [No]
-    4. (d)
-      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] Please refer to Section 4.1 for the detailed discussion.
-    5. (e)
-      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content?
-[No] This dataset is distributed across the highway, collecting traffic speed, traffic flow, and traffic occupancy.
-It does not contain personally identifiable information or offensive content.
-  5. 5.
-    If you used crowdsourcing or conducted research with human subjects…
-    1. (a)
-      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
-    2. (b)
-      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
-    3. (c)
-      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
 Appendix A Appendix
-A.1 Dataset
+A.1. Dataset
 To support our setting of extremely long forecasting with gaps, we selected a subset installed in the early stages and have consistently collected new data up to the present, which is shown as follows:
-Table 6: Overview of Gap, Hourly, and Daily Aggregated Data Based on Processed Raw Data
+Table 6. Overview of Gap, Hourly, and Daily Aggregated Data Based on Processed Raw Data
 Datasets(Gap/Hour/Day) Time Period Nodes
@@
 PEMS12_gap&agg 01/2002 - 03/2024 1543
-A.2 Results
+tfNSW 01/2013 - 05/2024 27
+A.2. Results
 Here we provided the results of District 5,6,10,11,12, as shown in Table 7 and Table 8:
-Table 7: Comparison in hourly and daily datasets in District 05,06,10,11,12. The symbol -- indicates that the result is an outlier.
+Table 7. Comparison in hourly and daily datasets in District 05,06,10,11,12. The symbol -- indicates that the result is an outlier.
 Methods Mamba iTransformer DLinear Autoformer
@@
 336 2.066 0.823 1.984 0.641 1.654 0.675 1.801 0.746
-Table 8: Comparison in gap dataset in District 05,06,10,11,12. The bold text indicates the best.
+Table 8. Comparison in gap dataset in District 05,06,07,10,11,12. The bold text indicates the best.
 Methods Mamba iTransformer DLinear Autoformer
@@
 336 0.945 0.710 1.550 0.970 1.415 0.934 0.955 0.739
+PEMS07_gap 1-year gap 96 1.719 1.006 1.756 1.024 1.680 1.024 1.215 0.867
+192 1.637 0.996 1.816 1.061 1.776 1.043 1.202 0.852
+336 1.637 0.991 1.784 1.028 1.774 1.034 1.098 0.774
+1.5-year gap 96 1.571 0.972 1.705 1.016 1.585 0.987 1.184 0.844
+192 1.578 0.982 1.657 1.004 1.624 0.996 1.088 0.800
+336 1.818 1.127 1.712 1.028 1.668 1.036 0.991 0.729
+2-year gap 96 5.411 1.319 2.055 1.149 1.746 1.048 1.099 0.802
+192 12.620 1.370 2.006 1.128 1.705 1.002 1.284 0.875
+336 9.614 1.378 1.812 1.057 1.605 0.968 0.972 0.734
 PEMS10_gap 1-year gap 96 2.310 1.233 0.882 0.698 1.466 0.963 1.208 0.865
@@
 336 0.814 0.628 0.789 0.631 0.776 0.624 0.754 0.617
-Additionally, we provided the results of naive zero-shot forecasting, as shown in Table 9. The poor performance of this method indicates significant potential for improvement.
-Table 9: Results of zero shot forecasting in PEMS03.
-In this study, we utilized the inputs from 151 existing nodes in PEMS03 to predict 52 new nodes, resulting in a test set comprising a total of 203 nodes.
-Datasets 96 192 336
-Metrics MSE MAE MSE MAE MSE MAE
-PEMS03_gap 1-year gap 4.658 3.465 4.945 3.648 5.198 3.892
-1.5-year gap 4.891 3.657 5.124 3.842 5.263 3.996
-2-year gap 4.547 3.410 4.895 3.539 5.103 3.758
+[…]
+Table 9. Model Training Time Comparison. The training time for all baselines per epoch is measured in seconds.
+Baselines Mamba iTransformer DLinear Autoformer Informer FEDFormer MICN PatchTST STGCN ASTGCN GWN AGCRN PDFormer
+PEMS04_gap1 396.3 1310.2 237.9 1656.7 652.5 900.9 409.8 3649.4 2425.9 7921.9 4110.4 13369.1 8013.9
+tfNSW_gap1 13.7 44.1 7.4 148.2 73.3 380.1 101.2 55.3 37.3 228.8 50.4 658.6 47.9

### v3

@@
 These limitations inevitably restrict the practical applicability of existing traffic forecasting datasets.
-To bridge this gap, we present XXLTraffic, largest available public traffic dataset with the longest timespan collected from Los Angeles, USA, and New South Wales, Australia, curated to support research in extremely long forecasting beyond test adaptation.
+To bridge this gap, we present XXLTraffic, the largest available public traffic dataset with the longest timespan collected from Los Angeles, USA, and New South Wales, Australia, curated to support research in extremely long forecasting beyond test adaptation.
 Our benchmark includes both typical time-series forecasting settings with hourly and daily aggregated data and novel configurations that introduce gaps and down-sample the training size to better simulate practical constraints.
@@
 Our dataset supplements existing spatio-temporal data resources and leads to new research directions in this domain.
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06copyright: acmlicensed
+Deep learning, Extremely long traffic forecasting
+copyright: acmlicensedjournalyear: 2025copyright: acmlicensedconference: The 33rd ACM International Conference on Advances in Geographic Information Systems; November 3–6, 2025; Minneapolis, MN, USAbooktitle: The 33rd ACM International Conference on Advances in Geographic Information Systems (SIGSPATIAL ’25), November 3–6, 2025, Minneapolis, MN, USAdoi: 10.1145/3748636.3762762isbn: 979-8-4007-2086-4/2025/11ccs: Information systems Sensor networks
 1. Introduction
@@
 In recent years, significant work has focused on both short-term and long-term traffic flow prediction.
-Deep learning techniques, including Graph Neural Networks (GNNs), have been employed to extract spatial relationships within traffic networks (Jin et al., 2023), while Transformer-based architectures have been utilized to capture temporal dependencies over various timescales (Shao et al., 2023a).
+Deep learning techniques, including Graph Neural Networks (GNNs), have been employed to extract spatial relationships within traffic networks (jin2023spatio), while Transformer-based architectures have been utilized to capture temporal dependencies over various timescales (shao2023exploring).
 Although these methods have shown promising results, they often rely on datasets that do not fully encapsulate the complexities introduced by rapid population growth and the surging number of vehicles, thus limiting their applicability to real-world scenarios.
 There is an emerging need in intelligent transportation systems to design predictive models that extend beyond test adaptation, effectively generalizing to real-world conditions that evolve over time.
-It is important to note that our concept of ’beyond test adaptation’ differs from ’test time adaptation’ (Guo et al., 2024) as illustrated in Fig.
-1 that shows the distinctions between them.
+It is important to note that our concept of ’beyond test adaptation’ differs from ’test time adaptation’  (guo2024online) as illustrated in Fig 1 that shows the distinctions between them.
 This shift requires models that can handle the multifaceted impacts of demographic changes and vehicle proliferation without relying solely on adaptation to specific test datasets.
@@
 Real-world traffic scenarios necessitate more complex prediction settings, involving extended temporal horizons or broader spatial coverage in experiments.
-In the temporal domain, new settings are typically proposed based on previously published work rather than introducing new datasets:  (Shao et al., 2023b) and (Jia et al., 2024) expanded input and output lengths up to four times on existing datasets.From a spatial perspective, (Chen et al., 2021) published a dataset with nodes growing annually and provided an evolving network to support new node predictions.
-(Wang et al., 2023b) proposed a continual learning framework with pattern expansion mechanisms based on (Chen et al., 2021).
-Additionally, SCPT (Prabowo et al., 2024) and Large-ST (Liu et al., 2024b) offered larger-scale spatial node datasets to support subsequent researchers.
+In the temporal domain, new settings are typically proposed based on previously published work rather than introducing new datasets:  (shao2023hutformer) and (jia2024witran) expanded input and output lengths up to four times on existing datasets.
+From a spatial perspective, (ijcai2021p0498) published a dataset with nodes growing annually and provided an evolving network to support new node predictions.
+(wang2023pattern) proposed a continual learning framework with pattern expansion mechanisms based on (ijcai2021p0498).
+Additionally, SCPT (prabowo2024traffic) and Large-ST (liu2024largest) offered larger-scale spatial node datasets to support subsequent researchers.
 Recent work has explored longer temporal step experimental settings and released traffic datasets spanning up to five years and thousands of nodes.
 However, in specific scenarios, such as future traffic prediction for highway planning, these data and experimental settings fall short.
-As shown in Fig.
-2, most existing datasets have limitations in temporal span, which inspired us to develop a dataset for expanding and extremely long traffic forecasting.
+As shown in Fig 2, most existing datasets have limitations in temporal span, which inspired us to develop a dataset for expanding and extremely long traffic forecasting.
 This need for traffic forecasting beyond test adaptation is crucial in various real-world scenarios.
@@
 1.2. Challenges
-The ultra-dynamic challenge encompasses three key aspects: (1) Continuously evolving states of the underlying spatio-temporal infrastructures, characterized by an expanding number of nodes over the years.
+The ultra-dynamic challenge encompasses two key aspects: (1) Continuously evolving states of the underlying spatio-temporal infrastructures, characterized by an expanding number of nodes over the years.
 This continuous growth introduces complexity as the infrastructure adapts and expands.
@@
 This dataset supports both temporally scalable and spatially scalable challenges in traffic prediction.
-  • We present an experimental setup with temporal gaps for extremely long prediction beyong test adaptation and provide a benchmark of aggregated versions of hourly and daily datasets.
+  • We present an experimental setup with temporal gaps for extremely long prediction beyond test adaptation and provide a benchmark of aggregated versions of hourly and daily datasets.
   • We provide exploration of input features through evolving temporal distributions over an extremely long observation period.
@@
 Traffic data primarily consists of vehicle flow detection data collected by sensors distributed across various locations in the traffic network.
-It is generally represented by $X_{i}$in $\mathbb{R}^{N\times T\times C}$, where $T$ denotes the time steps, $N$ denotes the number of sensors and $C$ denotes the number of features.
+It is generally represented by $X_{i}$ $\in$ $\mathbb{R}^{N\times T\times C}$, where $T$ denotes the time steps, $N$ denotes the number of sensors and $C$ denotes the number of features.
 Definition 2. Short-term traffic prediction:
@@
 As shown in Equation 1, the input length $\alpha$ and output length $\beta$ are generally set to 12 steps.
-(1) $[X_{t-(\alpha-1)},...,X_{t-1},X_{t}]\rightarrow[X_{t+1},X_{t+2},...,X_{t+\beta}],$
+(1) $X_{t-(\alpha-1)},...,X_{t-1},X_{t}]\rightarrow[X_{t+1},X_{t+2},...,X_{t+\beta}],$
 Definition 3. Long-term multivariate prediction:
@@
 Based on Equation 1, the observation and prediction are not adjacent but instead are separated by a gap period $g$, as shown in the following formula.
-(2) $[X_{t-(\alpha-1)},...,X_{t-1},X_{t}]\rightarrow[X_{t+g+1},X_{t+g+2},...,X_{t+g+\beta}],$
+(2) $X_{t-(\alpha-1)},...,X_{t-1},X_{t}]\rightarrow[X_{t+g+1},X_{t+g+2},...,X_{t+g+\beta}],$
 3. Gaps and Comparison with Existing Traffic Datasets
-Figure 3. Our expanding and extremely long prediction addresses the existing limitations in both short-term and long-term predictions.
 As shown in Table 1, existing traffic prediction work can easily be divided into short-term and long-term settings.
-The short-term setting originated from the STGCN(Yu et al., 2018) work, while the long-term setting was first introduced by LSTNet (Lai et al., 2018) and subsequently established as a widely adopted experimental framework by Informer (Zhou et al., 2021).
+The short-term setting originated from the STGCN(yu2018spatio) work, while the long-term setting was first introduced by LSTNet (lai2018modeling) and subsequently established as a widely adopted experimental framework by Informer (zhou2021informer).
 In recent years, short-term prediction typically has a maximum step length of 12 steps, while long-term prediction reaches up to 720 steps.
-However, works such as Witran (Jia et al., 2024) and DAN (Li et al., 2024) recognized the need for even longer step predictions in practical applications, extending the length to a maximum of four times the typical length.
+However, works such as Witran (jia2024witran) and DAN (Li_Xu_Anastasiu_2024) recognized the need for even longer step predictions in practical applications, extending the length to a maximum of four times the typical length.
 Despite the differences in step lengths, their observed and predicted values are concatenated tightly together, as shown in Equation 1.
 To accommodate complex real-world scenarios, such as highway route planning predictions, it is necessary to introduce a gap of several years between observation and prediction.
-Typically, existing datasets lack the temporal coverage required to support gaps exceeding one year.At the same time, predicting several years in advance also implies the need to forecast traffic for sensors at new locations, taking into account the evolving nature of the road network.
-Even when such coverage is available, works like (Wang et al., 2023b) and (Chen et al., 2021) utilize evolving datasets but do not provide sufficient data to train models for extended durations.
+Typically, existing datasets lack the temporal coverage required to support gaps exceeding one year.
+At the same time, predicting several years in advance also implies the need to forecast traffic for sensors at new locations, taking into account the evolving nature of the road network.
+Even when such coverage is available, works like (wang2023pattern) and (ijcai2021p0498) utilize evolving datasets but do not provide sufficient data to train models for extended durations.
 To overcome these gaps, our Expanding and Extremely Long Traffic Dataset robustly supports these complex scenarios.
@@
 Datasets Model Series Length
-Short-term STGCN (Yu et al., 2018) {3,6,9,12}
-DCRNN (Li et al., 2018), GWN (Wu et al., 2019), BTF (Chen and Sun, 2021), {3,6,12}
-DMSTGCN (Han et al., 2021), GTS (Shang et al., 2021), STGODE (Fang et al., 2021),
-PM-MemNet (Lee et al., 2021), STAEFormer (Liu et al., 2023)
-AGCRN (Bai et al., 2020), STSGCN (Song et al., 2020), ,DSTAGNN (Lan et al., 2022), {12}
-D2STGNN (Shao et al., 2022), DyHSL (Zhao et al., 2023),PDFormer (Jiang et al., 2023),
-MultiSPANS (Zou et al., 2024), GMSDR (Liu et al., 2022)
-Long-term MTGNN (Wu et al., 2020a),LSTNet (Lai et al., 2018) {3,6,12,24}
-ARU (Deshpande and Sarawagi, 2019) {12,24,48,168,336}
-LogSparse_Trans (Li et al., 2019) {24,48,72,96,120,144,168,192}
-AST (Wu et al., 2020b) {8,24,168,336}
-SSDNet (Lin et al., 2021) {20,24,30,138}
-Informer (Zhou et al., 2021), Autoformer (Wu et al., 2021), FEDformer (Zhou et al., 2022), {24,48,96,192,336,720}
-Linear (Li et al., 2022), Triformer (Cirstea et al., 2022), Pyraformer (Liu et al., 2021)
-DSformer (Yu et al., 2023),DeepTime (Woo et al., 2023),DLinear (Zeng et al., 2023)
-Witran (Jia et al., 2024) {168, 336, 720, 1440, 2880}
-DAN (Li et al., 2024) {288, 672, 1440}
-4. The XXLTraffic Datasets
+Short-term STGCN (yu2018spatio) {3,6,9,12}
+DCRNN (li2018diffusion), GWN (wu2019graph), BTF (chen2021bayesian), {3,6,12}
+DMSTGCN (han2021dynamic), GTS (shang2021discrete), STGODE (fang2021spatial),
+PM-MemNet (lee2021learning), STAEFormer (liu2023spatio)
+AGCRN (bai2020adaptive), STSGCN (song2020spatial), ,DSTAGNN (lan2022dstagnn), {12}
+D2STGNN (shao2022decoupled), DyHSL (zhao2023dynamic),PDFormer (jiang2023pdformer),
+MultiSPANS (zou2024multispans), GMSDR (liu2022msdr)
+Long-term MTGNN (wu2020connecting),LSTNet (lai2018modeling) {3,6,12,24}
+ARU (deshpande2019streaming) {12,24,48,168,336}
+LogSparse_Trans (li2019enhancing) {24,48,72,96,120,144,168,192}
+AST (wu2020adversarial) {8,24,168,336}
+SSDNet (lin2021ssdnet) {20,24,30,138}
+Informer (zhou2021informer), Autoformer (wu2021autoformer), FEDformer (zhou2022fedformer), {24,48,96,192,336,720}
+Linear (li2022simpler), Triformer (ijcai2022p277), Pyraformer (liu2021pyraformer)
+DSformer (yu2023dsformer),DeepTime (pmlr-v202-woo23b),DLinear (zeng2023transformers)
+Witran (jia2024witran) {168, 336, 720, 1440, 2880}
+DAN (Li_Xu_Anastasiu_2024) {288, 672, 1440}
+4. XXLTraffic Datasets
+(b) District 08 in 2005
+(c) District 08 in 2015
+(d) District 08 in 2024
+Figure 3.
+XXLTraffic dataset overview and its evolving development.
+This figure provides a global overview and two local overviews, showcasing the diversity of sensor distribution.
+The lower part highlights a selected region to illustrate the growth and changes in traffic sensors over time.
 4.1. Data Collection
-We obtained the expanding and extremely long traffic sensor data from the California Department of Transportation (CalTrans) Performance Measurement System [^1] https://pems.dot.ca.gov/ (PeMS) (Chen et al., 2001) and Transport for NSW [^2] https://maps.transport.nsw.gov.au/egeomaps/traffic-volumes/index.html.
+We obtained the expanding and extremely long traffic sensor data from the California Department of Transportation (CalTrans) Performance Measurement System [^1] https://pems.dot.ca.gov/ (PeMS) (chen2001freeway) and Transport for NSW [^2] https://maps.transport.nsw.gov.au/egeomaps/traffic-volumes/index.html.
 PeMS is an online platform that collects traffic data from 19,561 sensors distributed across California state highways.
@@
 We have stored the complete raw data files in an open-source repository for quick access, which will be released after the publication.
-The tfNSW is an open-source data platform provided by Transport for NSW, featuring traffic flow data collected from sensors distributed along major roads throughout the state of New South Wales of Australia.
+TfNSW is an open-source data platform provided by Transport for NSW, featuring traffic flow data collected from sensors distributed along major roads throughout the state of New South Wales of Australia.
 The data is available at a minimum granularity of one hour.
+As TfNSW’s data is integrated into a visualization webpage, it can only be retrieved via manual, item-by-item downloads, after which independent aggregation and preprocessing are required.
 Table 2. Comparison between our XXLTraffic dataset and the existing traffic datasets.
@@
 As illustrated in Table 2, our collected dataset significantly exceeds existing datasets in terms of both temporal coverage and the number of spatial nodes.
-The dataset sample will be available on: https://anonymous.4open.science/r/XXLTraffic-F281, which includes the raw data, sensor metadata (containing sensor IDs, geographical coordinates, associated road information, etc.), the data processing pipeline code, and the processed datasets.
+The dataset sample will be available on: https://github.com/cruiseresearchgroup/XXLTraffic, which includes the raw data, sensor metadata (containing sensor IDs, geographical coordinates, associated road information, etc.), the data processing pipeline code, and the processed datasets.
 4.2. Data Preprocessing
 Based on the 23 years of raw data we collected, we conducted rigorous data filtering and aggregation.
-The PeMS system has continuously evolved, expanding from a few sensors in 2001 to over 4,000 sensors in some districts today.
-To support our setting of extremely long forecasting with gaps, we selected a subset of sensors that were installed in the early stages and have consistently collected new data up to the present(named gap dataset), which is shown in the Appendix.
-This extensive gap dataset effectively underpins the extremely long forecasting with gaps demonstrated in Fig.
-3.
+PeMS system has continuously evolved, expanding from a few sensors in 2001 to over 4,000 sensors in some districts today.
+To support our setting of extremely long forecasting with gaps, we selected a subset of sensors that were installed in the early stages and have consistently collected new data up to the present (named gap dataset), which is shown in Table 5.
+This extensive gap dataset effectively underpins the extremely long forecasting with gaps demonstrated in Fig 5.
 Utilizing the gap dataset, we performed both hourly and daily aggregations, which will be employed for gap-free long-term forecasting benchmarking.
 We will provide standard long-term forecasting benchmarks for both the hourly and daily datasets.
-(b) District 08 in 2005
-(c) District 08 in 2015
-(d) District 08 in 2024
-Figure 4.
-XXLTraffic dataset overview and its evolving development.
-This figure provides a global overview and two local overviews, showcasing the diversity of sensor distribution.
-The lower parts highlights a selected region to illustrate the growth and changes in traffic sensors over time.
+The data preprocessing code has also been made available in the GitHub repository.
 4.3. Data Overview
-The XXLTraffic dataset is distributed across highways in the state of California, as illustrated in the Fig.
-4(a).
-The nine colors represent nine districts.
-From Figures.
-4(b),  4(c), and 4(d), we can clearly observe the evolutionary growth of the sensors.
+XXLTraffic dataset is distributed across highways in the state of California in USA and the traffic flow in NSW state in Australia, as illustrated in the Fig 3(a).
+From Figures 3(b),  3(c), and 3(d), we can clearly observe the evolutionary growth of the sensors.
 The sensors are extensively distributed across both urban and suburban areas, offering diverse modalities.
 Additionally, the sensors are densely interconnected, enabling the formation of a high-quality traffic graph dataset.
-It is evident that sensors at the same location may collect completely different distributions over the course of urban evolution.
-As shown in Fig.
-5, some sensors have maintained the same distribution from 2005 to 2024, while others have experienced significant changes in distribution.
-The temporal changes causing domain shifts present a significant challenge for our extremely long forecasting.
 (a) distribution of District 8 in PeMS.
@@
 Figure 4.
-Sensor traffic status distribution of District 8 in PeMS from 2005 to 2024 in 5(a) and from 2016 to 2022 in NSW in 5(b).
+Sensor traffic status distribution of District 8 in PeMS from 2005 to 2024 in 4(a) and from 2016 to 2022 in NSW in 4(b).
 While some sensors exhibit minimal changes, others show significant distribution differences, regardless of whether they are in low-traffic or high-traffic areas.
 This presents substantial challenges for extremely long forecasting with long gaps.
+It is evident that sensors at the same location may collect completely different distributions over the course of urban evolution.
+As shown in Fig 4, which depicts the kernel density plots of sensor values, some sensors have maintained the same distribution from 2005 to 2024, while others have experienced significant changes in distribution.
+Temporal changes causing domain shifts present a significant challenge for our extremely long forecasting.
 4.4. XXLTraffic Licence
-The XXLTraffic dataset is licensed under CC BY-NC 4.0 International: https://creativecommons.org/licenses/by-nc/4.0.
+XXLTraffic dataset is licensed under CC BY-NC 4.0 International: https://creativecommons.org/licenses/by-nc/4.0.
 Our code is available under the MIT License: https://opensource.org/licenses/MIT.
@@
 We conducted experiments for both extremely long forecasting with gaps using gap dataset and conventional long-term forecasting using hourly dataset and daily dataset.
-Additionally, referring to the definition in Fig.
-3, we set the gap parameter $g$ as 1 year, 1.5 years, and 2 years for the gap dataset, as illustrated by Fig.
-6.
+Additionally, referring to the definition in Fig 5, we set the gap parameter $g$ as 1 year, 1.5 years, and 2 years for the gap dataset, as illustrated by Fig 5.
 Figure 5. Problem definition. The yellow boxes represent typical predictions, the gray boxes denote gap periods between observation and prediction, and the blue boxes indicate extended predictions.
@@
 We conducted experiments on all proposed sub-datasets.
-To maintain consistency with previous state-of-the-art benchmarks, we selected districts 03, 04, and 08 (widely recognized as PEMS03/04/08) for the experiments using the gap, and districts 03, 04,07 and 08 for hourly, and daily experiments.
-Results for other datasets are presented in Appendix.
+To maintain consistency with previous state-of-the-art (SOTA) benchmarks, we selected districts 03, 04, and 08 (widely recognized as PEMS03/04/08) for the experiments using the gap which is shown in Table 3.
+Results for other gaps and aggregated datasets on some baselines are presented in Tables˜4, 8 and 9.
 All sub-datasets were divided into training, validation, and test sets using a 6:2:2 ratio.
 For the gap dataset, due to the extensive span of up to 20 years resulting in a large sample size, we fixed a seed during data preprocessing to select 10% of the dataset for training and testing to quickly demonstrate our results.
-The details of the datasets used in our benchmarking is in Appendix A.1.
-Table 3. Comparison in gap dataset. The bold text indicates the best.
+Details of the datasets used in our benchmarking is in Table 5.
+Table 3. Comparison in gap dataset. Bold numbers indicate the best results.
 Gap Data Gap Metric Horizon Mamba iTrans DLinear Autofor Infor FEDFor MICN PatchTST STGCN ASTGCN GWN AGCRN PDFor
@@
 336 0.903 0.730 0.748 0.849 0.639 0.766 0.691 0.714 0.810 0.693 0.695 0.749 0.715
+Table 4. Comparison in datasets in District 05,06,07,10,11,12. Bold numbers indicate the best results.
+Datasets PEMS05 PEMS06
+Metrics 1-year 1.5-year 2-year 1-year 1.5-year 2-year
+Mamba 96 MSE/MAE 2.079 / 1.209 1.852 / 1.122 1.868 / 1.106 1.806 / 1.066 1.549 / 0.997 1.226 / 0.851
+192 MSE/MAE 2.132 / 1.256 1.929 / 1.182 2.219 / 1.274 1.928 / 1.112 1.746 / 1.054 0.949 / 0.720
+336 MSE/MAE 2.377 / 1.340 2.370 / 1.313 2.695 / 1.212 2.181 / 1.212 1.605 / 1.018 0.945 / 0.710
+iTransformer 96 MSE/MAE 1.945 / 1.164 1.879 / 1.078 1.580 / 0.969 0.875 / 0.692 1.331 / 0.891 1.864 / 1.106
+192 MSE/MAE 1.984 / 1.185 1.593 / 1.032 1.481 / 0.958 1.227 / 0.848 1.077 / 0.778 1.343 / 0.879
+336 MSE/MAE 2.067 / 1.234 2.214 / 1.071 2.207 / 1.201 1.594 / 1.003 1.500 / 0.961 1.550 / 0.970
+DLinear 96 MSE/MAE 1.291 / 0.916 1.683 / 1.054 1.602 / 1.018 1.173 / 0.837 1.484 / 0.963 1.691 / 1.033
+192 MSE/MAE 1.750 / 1.099 1.633 / 1.045 1.589 / 1.027 1.410 / 0.942 1.353 / 0.920 1.259 / 0.858
+336 MSE/MAE 1.894 / 1.144 0.794 / 1.184 1.922 / 1.139 1.501 / 0.976 1.587 / 1.011 1.415 / 0.934
+Autoformer 96 MSE/MAE 1.065 / 0.796 1.060 / 0.785 0.828 / 0.672 1.216 / 0.859 0.885 / 0.710 1.013 / 0.768
+192 MSE/MAE 1.063 / 0.809 0.912 / 0.712 1.018 / 0.772 0.961 / 0.751 1.010 / 0.768 0.853 / 0.691
+336 MSE/MAE 1.135 / 0.827 1.345 / 0.882 1.186 / 0.839 0.992 / 0.769 0.955 / 0.739 0.955 / 0.739
+Table 5. Overview of Gap, Hourly, and Daily aggregated data based on processed raw data
+Datasets(Gap/Hour/Day) Time Period Nodes
+PEMS03_gap&agg 03/2001 - 03/2024 151
+PEMS04_gap&agg 06/2002 - 03/2024 822
+PEMS05_gap&agg 03/2012 - 03/2024 103
+PEMS06_gap&agg 12/2009 - 03/2024 130
+PEMS07_gap&agg 06/2002 - 03/2024 3062
+PEMS08_gap&agg 03/2001 - 03/2024 212
+PEMS10_gap&agg 06/2007 - 03/2024 107
+PEMS11_gap&agg 09/2002 - 03/2024 521
+PEMS12_gap&agg 01/2002 - 03/2024 1543
+tfNSW 01/2013 - 05/2024 27
 5.2. Baselines
 In our comparison experiments, we adopted four popular baselines, including MLP, Transformer, and Mamba architectures.
-Informer (Zhou et al., 2021) introduces an efficient transformer for long sequence time-series forecasting using ProbSparse self-attention and self-attention distilling, enabling encoder-decoder architectures to handle long sequences effectively.
-MICN (Wang et al., 2023a) proposes a multi-scale context network that models both local and global contexts for long-term time series forecasting, capturing patterns across different temporal scales to enhance performance.
-FEDformer (Zhou et al., 2022) introduces a frequency-enhanced decomposed transformer that models time series in both time and frequency domains, improving long-term forecasting by effectively capturing temporal patterns.
-PatchTST (Nie et al., [n.
-d.]) applies transformers to time series by treating them as sequences of patches, enabling effective long-term forecasting through self-attention over patch representations to capture temporal dependencies.Autoformer (Wu et al., 2021), an earlier state-of-the-art model, leverages a decomposition architecture and auto-correlation mechanism to enhance efficiency and accuracy in long-term time series forecasting, outperforming traditional Transformer models.
-iTransformer (Liu et al., 2024a) is the latest and most effective Transformer-based model, utilizing attention and feed-forward networks on inverted dimensions, embedding time points into variate tokens.
-DLinear (Zeng et al., 2023) challenges the effectiveness of Transformer models by proposing a simple one-layer linear model that captures temporal relations in an ordered set of continuous points.
+Informer (zhou2021informer) introduces an efficient transformer for long sequence time-series forecasting using ProbSparse self-attention and self-attention distilling, enabling encoder-decoder architectures to handle long sequences effectively.
+MICN (wang2023micn) proposes a multi-scale context network that models both local and global contexts for long-term time series forecasting, capturing patterns across different temporal scales to enhance performance.
+FEDformer (zhou2022fedformer) introduces a frequency-enhanced decomposed transformer that models time series in both time and frequency domains, improving long-term forecasting by effectively capturing temporal patterns.
+PatchTST (nietime) applies transformers to time series by treating them as sequences of patches, enabling effective long-term forecasting through self-attention over patch representations to capture temporal dependencies.Autoformer (wu2021autoformer), an earlier SOTA model, leverages a decomposition architecture and auto-correlation mechanism to enhance efficiency and accuracy in long-term time series forecasting, outperforming traditional Transformer models.
+iTransformer (liu2024itransformer) is the latest and most effective Transformer-based model, utilizing attention and feed-forward networks on inverted dimensions, embedding time points into variate tokens.
+DLinear (zeng2023transformers) challenges the effectiveness of Transformer models by proposing a simple one-layer linear model that captures temporal relations in an ordered set of continuous points.
 It employs positional encoding and uses tokens to embed sub-series, preserving some ordering information in Transformers.
-Lastly, Mamba (Gu and Dao, 2023), a well-known sequential model from last year, uses a bidirectional Mamba block to extract inter-variate correlations and temporal dependencies.
-Additionally, we have selected five SOTA baselines  (Yu et al., 2018; Guo et al., 2019; Wu et al., 2019; Bai et al., 2020; Jiang et al., 2023) from traffic forecasting domain.
+Lastly, Mamba (gu2023mamba), a well-known sequential model from last year, uses a bidirectional Mamba block to extract inter-variate correlations and temporal dependencies.
+Additionally, we have selected five SOTA baselines  (yu2018spatio; guo2019attention; wu2019graph; bai2020adaptive; jiang2023pdformer) from traffic forecasting domain.
+Meanwhile, as shown in Table 6, whether using 10% of the data, 50% of the data, or the full dataset, early stopping still occurred within the first or second epoch, and the results remained largely similar.
+This suggests that the amount of sampled training data does not substantially affect the outcomes, as the task itself is inherently difficult.
+This further highlights the major challenge introduced in this work: how to effectively leverage the available data to forecast traffic flow far into the future.
 5.3. Implementation Details
-We adopted the default settings of the Time-Series-Library (Wu et al., 2022) to conduct a comprehensive comparison of baselines.
+We adopted the default settings of the Time-Series-Library (wu2022timesnet) to conduct a comprehensive comparison of baselines.
 We use the results of five random seeds as the average.
-We used 96 time steps as input and 336 time steps as ground truth.
+We used 96 time steps as input to predict 96, 192, 336 steps.
 The code was implemented in PyTorch and executed on a V100 GPU with 32GB memory and 384GB RAM, provided by NCI Australia, an NCRIS-enabled capability supported by the Australian Government.
+Table 6. Results of different fractions of data in MICN.
+Fraction of Data 10% 50% 100%
+Metrics MSE MAE MSE MAE MSE MAE
+PEMS03_gap 1-year 96 0.514 0.515 0.990 0.736 0.910 0.683
+192 0.528 0.513 0.889 0.672 0.814 0.651
+336 0.493 0.491 0.808 0.650 0.744 0.612
 5.4. Results of Extremely Long Forecasting with Gaps
@@
 It is observed that nearly all results are poor, highlighting the significant challenge posed by domain shifts over time for extremely long forecasting with gaps.
-These baseline results also indicate that traditional the-state-of-the-art (SOTA) rankings and methodologies are no longer effective.
+These baseline results also indicate that traditional SOTA rankings and methodologies are no longer effective.
 Notably, MICN, which performs the worst under conventional data and settings, shows the best performance in our setting.
@@
 This insight suggests that future efforts to tackle this problem should place greater emphasis on leveraging the intrinsic potential of the data.
-Additionally, we have compiled the training time for one epoch of all the baselines on the largest dataset, PEMS04_gap, and the smallest dataset, tfNSW, as shown in Table 9.
+Additionally, we have compiled the training time for one epoch of all the baselines on the largest dataset, PEMS04_gap, and the smallest dataset, tfNSW, as shown in Figure 6.
+The reason the PEMS08 dataset yields missing results in Informer is that it is completely unfit for modeling, resulting in extremely poor performance.
 Considering that our dataset provides an extensive temporal range for training, we can theoretically extend the input step length considerably.
@@
 As shown in Table 7, the performance improves significantly with the increase in input step length, further demonstrating the substantial potential of our dataset for deep exploration.
+And the results of datasets PEMS05_gap, PEMS06_gap, PEMS07_gap, PEMS10_gap, PEMS11_gap, PEMS12_gap are shown in Table 4.
 Table 7. Results of ablation study with 4 different input step lengths
@@
 5.5. Results of Hourly and Daily Forecasting
-We believe that both the hourly and daily datasets are equally significant.
+As shown in Table 8, we believe that both the hourly and daily datasets are equally significant.
 Multi-scale, diverse datasets can provide the community with valuable references.
@@
 Smaller time scales help reduce complexity and uncertainty, thereby improving the accuracy of prediction.
-Research has shown that clustering at different scales can improve model performance (Wang et al., 2024).
+Research has shown that clustering at different scales can improve model performance (wang2024timemixer).
 Therefore, our aggregated version of the data will contribute new external features to the community.
@@
 336 1.161 0.697 1.024 0.617 1.022 0.602 1.127 0.689
-6. Prospects and Constraints
-Prospects.
-Our dataset spans the longest time period among existing datasets, and it is not only the largest spatially, but also evolving in growth.
-It can continue to update according to the updates from the PeMS system in the future.
-It is specifically designed for various complex scenarios, such as those already mentioned, including extremely long forecasting with long gaps, and hourly and daily predictions.
-Constraints.
-The limitations of our dataset are also quite evident.
-Due to the sheer size of our dataset, it requires more computational resources.
-However, with the maturation of large language models and foundational models, we believe its large volume will become an advantage, contributing more diverse data to the spatio-temporal large model community.
-Appendix A Appendix
-A.1. Dataset
-To support our setting of extremely long forecasting with gaps, we selected a subset installed in the early stages and have consistently collected new data up to the present, which is shown as follows:
-Table 6. Overview of Gap, Hourly, and Daily Aggregated Data Based on Processed Raw Data
-Datasets(Gap/Hour/Day) Time Period Nodes
-PEMS03_gap&agg 03/2001 - 03/2024 151
-PEMS04_gap&agg 06/2002 - 03/2024 822
-PEMS05_gap&agg 03/2012 - 03/2024 103
-PEMS06_gap&agg 12/2009 - 03/2024 130
-PEMS07_gap&agg 06/2002 - 03/2024 3062
-PEMS08_gap&agg 03/2001 - 03/2024 212
-PEMS10_gap&agg 06/2007 - 03/2024 107
-PEMS11_gap&agg 09/2002 - 03/2024 521
-PEMS12_gap&agg 01/2002 - 03/2024 1543
-tfNSW 01/2013 - 05/2024 27
-A.2. Results
-Here we provided the results of District 5,6,10,11,12, as shown in Table 7 and Table 8:
-Table 7. Comparison in hourly and daily datasets in District 05,06,10,11,12. The symbol -- indicates that the result is an outlier.
+Here we also provided the results of District 5,6,10,11,12, as shown in Table 9 and Table 4.
+The reason why some datasets yield missing results is that they are completely unfit for modeling, resulting in extremely poor performance.
+Table 9. Comparison in hourly and daily datasets in District 05,06,10,11,12. The symbol - indicates that the result is an outlier.
 Methods Mamba iTransformer DLinear Autoformer
@@
 336 2.066 0.823 1.984 0.641 1.654 0.675 1.801 0.746
-Table 8. Comparison in gap dataset in District 05,06,07,10,11,12. The bold text indicates the best.
-Methods Mamba iTransformer DLinear Autoformer
-Metrics MSE MAE MSE MAE MSE MAE MSE MAE
-PEMS05_gap 1-year gap 96 2.079 1.209 1.945 1.164 1.291 0.916 1.065 0.796
-192 2.132 1.256 1.984 1.185 1.750 1.099 1.063 0.809
-336 2.377 1.340 2.067 1.234 1.894 1.144 1.135 0.827
-1.5-year gap 96 1.852 1.122 1.879 1.078 1.683 1.054 1.060 0.785
-192 1.929 1.182 1.593 1.032 1.633 1.045 0.912 0.712
-336 2.370 1.313 2.214 1.071 0.794 1.184 1.345 0.882
-2-year gap 96 1.868 1.106 1.580 0.969 1.602 1.018 0.828 0.672
-192 2.219 1.274 1.481 0.958 1.589 1.027 1.018 0.772
-336 2.695 1.212 2.207 1.201 1.922 1.139 1.186 0.839
-PEMS06_gap 1-year gap 96 1.806 1.066 0.875 0.692 1.173 0.837 1.216 0.859
-192 1.928 1.112 1.227 0.848 1.410 0.942 0.961 0.751
-336 2.181 1.212 1.594 1.003 1.501 0.976 0.992 0.769
-1.5-year gap 96 1.549 0.997 1.331 0.891 1.484 0.963 0.885 0.710
-192 1.746 1.054 1.077 0.778 1.353 0.920 1.010 0.768
-336 1.605 1.018 1.500 0.961 1.587 1.011 0.955 0.739
-2-year gap 96 1.226 0.851 1.864 1.106 1.691 1.033 1.013 0.768
-192 0.949 0.720 1.343 0.879 1.259 0.858 0.853 0.691
-336 0.945 0.710 1.550 0.970 1.415 0.934 0.955 0.739
-PEMS07_gap 1-year gap 96 1.719 1.006 1.756 1.024 1.680 1.024 1.215 0.867
-192 1.637 0.996 1.816 1.061 1.776 1.043 1.202 0.852
-336 1.637 0.991 1.784 1.028 1.774 1.034 1.098 0.774
-1.5-year gap 96 1.571 0.972 1.705 1.016 1.585 0.987 1.184 0.844
-192 1.578 0.982 1.657 1.004 1.624 0.996 1.088 0.800
-336 1.818 1.127 1.712 1.028 1.668 1.036 0.991 0.729
-2-year gap 96 5.411 1.319 2.055 1.149 1.746 1.048 1.099 0.802
-192 12.620 1.370 2.006 1.128 1.705 1.002 1.284 0.875
-336 9.614 1.378 1.812 1.057 1.605 0.968 0.972 0.734
-PEMS10_gap 1-year gap 96 2.310 1.233 0.882 0.698 1.466 0.963 1.208 0.865
-192 2.878 1.396 1.606 0.993 1.834 1.106 1.152 0.860
-336 2.584 1.327 1.413 0.825 1.887 1.121 1.130 0.830
-1.5-year gap 96 2.181 1.206 1.812 1.052 1.765 1.057 1.368 0.913
-192 2.525 1.318 1.645 0.986 1.791 1.069 1.151 0.828
-336 2.488 1.310 2.051 1.134 1.976 1.139 1.111 0.809
-2-year gap 96 1.165 0.822 2.772 1.383 1.977 1.119 1.188 0.845
-192 1.082 0.792 1.974 1.084 1.490 0.932 0.971 0.751
-336 1.021 0.764 2.157 1.168 1.619 0.989 1.096 0.809
-PEMS11_gap 1-year gap 96 5.199 0.891 5.417 0.997 5.276 0.936 4.930 0.854
-192 5.300 0.920 5.504 1.029 5.393 0.974 5.114 0.867
-336 5.251 0.901 5.410 0.985 5.364 0.955 4.830 0.849
-1.5-year gap 96 5.871 0.968 6.045 1.296 5.914 1.224 5.691 0.878
-192 5.968 1.012 6.121 1.314 5.993 1.229 5.792 0.884
-336 6.167 1.074 6.214 1.326 6.025 1.299 5.947 0.967
-2-year gap 96 5.914 0.996 6.136 1.318 5.945 1.243 5.761 0.978
-192 6.541 1.043 6.213 1.327 6.014 1.289 6.245 1.001
-336 6.541 1.086 6.221 1.332 6.024 1.300 6.268 1.024
-PEMS12_gap 1-year gap 96 1.751 1.025 1.624 1.002 1.611 1.005 1.060 0.789
-192 1.726 1.024 1.424 0.929 1.537 0.972 1.150 0.840
-336 1.751 1.029 1.672 1.015 1.683 1.017 0.889 0.719
-1.5-year gap 96 1.554 0.967 1.479 0.921 1.468 0.910 0.954 0.875
-192 1.401 0.898 1.314 0.867 1.301 0.862 0.943 0.869
-336 1.417 0.906 1.322 0.869 1.298 0.859 0.921 0.846
-2-year gap 96 0.956 0.704 0.876 0.671 0.872 0.664 0.846 0.659
-192 0.846 0.656 0.813 0.653 0.806 0.649 0.785 0.628
-336 0.814 0.628 0.789 0.631 0.776 0.624 0.754 0.617
-[…]
-Table 9. Model Training Time Comparison. The training time for all baselines per epoch is measured in seconds.
-Baselines Mamba iTransformer DLinear Autoformer Informer FEDFormer MICN PatchTST STGCN ASTGCN GWN AGCRN PDFormer
-PEMS04_gap1 396.3 1310.2 237.9 1656.7 652.5 900.9 409.8 3649.4 2425.9 7921.9 4110.4 13369.1 8013.9
-tfNSW_gap1 13.7 44.1 7.4 148.2 73.3 380.1 101.2 55.3 37.3 228.8 50.4 658.6 47.9
+5.6. Training Efficiency Analysis
+We have provided Figure 6 that compares training times per epoch (in seconds) for various models on PEMS04_gap1 and tfNSW_gap1 datasets.
+DLinear is the fastest $237.9\text{s}$ and $7.4\text{s}$, while AGCRN is the slowest $13{,}369.1\text{s}$ and $658.6\text{s}$.
+Mamba shows competitive efficiency $396.3\text{s}$ and $13.7\text{s}$, while PatchTST and FEDFormer have long training times.
+Overall, DLinear is the most efficient, while AGCRN and PDFormer are computationally intensive.
+Figure 6. Model training time comparison. The training time for all baselines per epoch is measured in seconds.
+6. Conclusion
+In this work, we introduced XXLTraffic, the largest and longest‐spanning public traffic dataset to date, covering over 23 years of sensor data from both Los Angeles (PeMS) and New South Wales (TfNSW).
+By selecting consistently active sensors and providing both hourly and daily aggregations, we deliver a versatile resource that reflects continuous spatial growth and evolving temporal distributions.
+To support forecasting “beyond test adaptation,” we curated specialized gap‐based subsets (with 1, 1.5 and 2 year intervals) and established benchmark configurations that expose models to real‐world domain shifts, alongside traditional short- and long-term prediction settings.
+Our dataset release—including raw data, metadata, preprocessing code, and processed aggregates—offers the community a robust platform for developing and evaluating models under a range of practical constraints and forecasting horizons.
+While the unprecedented scale and richness of XXLTraffic unlocks new research directions in spatio-temporal modeling and large-model pretraining, its sheer size does impose significant computational demands.
+Future work should explore efficient data-subsampling strategies, scalable training algorithms, and integration with emerging large-language and foundation models to fully leverage the dataset’s potential without prohibitive resource requirements.
+7. Acknowledgments
+We would like to acknowledge the support of Cisco’s National Industry Innovation Network (NIIN) Research Chair Program and the ARC Centre of Excellence for Automated Decision-Making and Society (CE200100005).
+We acknowledge the resources and services from the National Computational Infrastructure (NCI), which is supported by the Australian Government.